### v1

# Simple Network Mechanism Leads to Quasi-Real Brain Activation Patterns with Drosophila Connectome

Abstract
Considering the high computational demands of most methods, using network communication models to simulate the brain is a more economical way. However, despite numerous brain network communication models, there is still insufficient evidence that they can effectively replicate the real activation patterns of the brain. Moreover, it remains unclear whether actual network structures are crucial in simulating intelligence. Addressing these issues, we propose a large scale network communication model based on simple rules and design criteria to assess the differences between network models and real situations. We conduct research on the biggest adult Drosophila connectome data set. Experimental results show significant activation in neurons that should respond to stimulus and slight activation in irrelevant ones, which we call quasi-real activation pattern. Besides, when we change the network structure, the quasi-activation patterns disappear. Interestingly, activation regions have shorter network distances to their input neurons, implying that the network structure (not spatial distance) is the core to form brain functionality. In addition, giving the input neurons a unilateral stimulus, we observe a bilateral response, which is consistent with reality. Then we find that both hemispheres have extremely similar statistical indicators. We also develop real-time 3D large spatial network visualization software to observe and document experimental phenomena, filling the software gap. This research reveals network models power: it can reach the quasi-activation pattern even with simple propagation rules. Besides, it provides evidence that network structure matters in brain activity pattern generation. Future research could fully simulate brain behavior through network models, paving the way for artificial intelligence by developing new propagation rules and optimizing link weights.
Keywords: brain network, network communication model, drosophila connectome, activation pattern
1 Introduction
The neural systems of intelligent beings maintain constant communication[1][2]. The intricate mechanism of this communication unfolds as follows: specific neurons, often input neurons, receive stimulus signals from the external environment or other parts of the body, transitioning from a resting to an activated state. Subsequently, they relay these stimulus signals to the next tier of neurons through synapses. This process gives rise to macroscopic emergence across the entire neural system. Such communication fosters the emergence of intelligence in all intelligent life forms, including humans. Consequently, comprehending and replicating this process has become an indispensable avenue for unraveling biological intelligence and realizing artificial intelligence[3].
Neural systems are interconnected networks[4], with neurons engaging in constant communication via synapses within the network’s topology. The brain’s unique and complex network structure is pivotal to the emergence of intelligence[5]. Many scientists posit that network science is crucial for deciphering the mechanisms underlying the flexible regulation of neural communication[6][7][8]. Determining how to leverage network science methodologies stands as a paramount challenge in contemporary neuroscience. Within this discipline, neural networks are often represented as graphs, consisting of nodes (entities) and edges (relationships). From the broadcasting model[9] to the diffusion model[10], researchers strive to craft network models that mimic activation patterns observed in the brain, offering insights into the workings of neural networks.
However, despite artificial neural networks (ANNs)’ deviation from the brain’s actual structure, they get full achievements in achieving intelligence, which cast doubt on the necessity of real network structures in brain behavior simulation. Furthermore, some researchers have highlighted the absence of evidence linking network model simulations to genuine brain activity patterns[11], undermining the models’ status as a robust research tool in neuroscience. The effectiveness of network models in replicating the neuronal system’s activation dynamics remains uncertain. This uncertainty partly stems from many models’ failure to replicate the brain’s structure at the neuron and synapse level, compromising the models’ interpretability.
Moreover, current brain network communication models often rely on prevalent metrics such as signaling cost and activation time to assess model quality[11]. Yet, these metrics fall short of verifying whether a network model accurately reflects the brain’s actual activation patterns, which is the fundamental goal of brain simulation. There is a need for evaluation criteria that can affirm whether network communication models genuinely mirror the brain’s communication processes. Additionally, observing the brain network communication process is important for researches. The absence of 3D visualization software capable of real-time monitoring of communication status and displaying networks with a vast number of nodes poses challenges for researchers, including neuroscientists, by impeding real-time observation, display, and analysis of experimental outcomes.
Over the past few decades, biologists have dedicated themselves to determining the precise structure of the connectome, ranging from C. elegans[12] to Drosophila larvae[13]. By understanding the connections within an organism’s nervous system, researchers can gain insights into how the nervous system operates. A recent study[14] has published the largest adult Drosophila connectome, providing data support for researches in network science.
For these reasons, we conducted experiments utilizing the recently released comprehensive data set of the largest adult Drosophila connectome, enhancing the interpretability of our findings. Our approach involves crafting a large scale network communication model based on simple mechanisms, designed to simulate the activation behavior observed in the connectome.
Specifically, We establish the network representation of the connectome. Then we design each neuron’s state and threshold functions. When a neuron’s state meets the threshold, it activates and subsequently transmits signals to other neurons. After a few iteration, the whole network emerges an activation pattern. Additionally, we propose an evaluation criteria based on the ratio of actually activated neurons to the expected number of activation. Experiments reveals that our model has significant activation in the type of neurons that should be affected by the stimulus and slight activation in irrelevant neurons which called a quasi-real activation pattern. To prove the vitalness of network structure, we calculate every node’s layer and shuffle part of the network structure in same layer and find the shuffled network can’t generate any quasi-real activation pattern.
To enhance our understanding, we developed a 3D network visualization software, utilizing HTML5’s Canvas and Three.js based on WebGL technology for large spatial network visualization effects. This tool enables real-time observation of each neuron’s communication process.
To prove the network structure matters (rather than spatial structure) in activation pattern generation, we calculate the average network and spatial distance of input and activation neurons and results reveal that the average network distance between input neurons and their corresponding activation areas is shorter than unrelated areas, which is a strong evidence to prove network structure matters in the process of message transmission in brain.
Furthermore, by subjecting input neurons to the same stimulus, we observe significant responses in neurons expected to be affected, with only slight responses in irrelevant neurons—characterized as a quasi-real activation pattern. Notably, when providing unilateral stimuli to input neurons, we observe a bilateral response in the Drosophila neuronal wiring diagram, aligning with reality. Then we calculate the statistics properties and find that the left and right hemispheres have extremely similar degree, clustering coefficient, and eigenvector centrality vector coefficient. The vectors’ Person coefficient are up to 0.9986, 0.9988, and 0.9991, respectively.
2 Related Works
Various models have been developed to mimic or understand aspects of brain function, ranging from the cellular level to the whole brain. In this section we review the models that aim to simulate the brain’s behavior. The review includes neural dynamic model that describes how action potentials in neurons are initiated and propagated. In addition, we introduce the large scale brain simulating methods. At last, we introduce the network communication models which refers to a computational or theoretical framework designed to understand and simulate how different parts of the brain communicate with each other.
2.1 Neuron Dynamic Model
In past years, neuroscientists have aimed to describing the behavior of neurons by dynamic model. One of the most famous works is the Hodgkin-Huxley (HH) model[15]. It includes a set of differential equations that describe the changes in membrane potential and the dynamics of ion channels. The key variables in the model are the membrane potential $V$, and the gating variables $m$, $n$, $h$ that represent the probability of ion channels being open.
The currents are given by:
$I_{Na}$ $=\bar{g}{Na}m^{3}h(V-E{Na})$ (1)
$I_{K}$ $=\bar{g}_{K}n^{4}(V-E_{K})$ (2)
$I_{L}$ $=\bar{g}_{L}(V-E_{L})$ (3)
Where $\bar{g}_{Na}$, $\bar{g}_{K}$ and $\bar{g}_{L}$ are the maximum conductances for sodium, potassium, and leak channels. $E_{Na}$, $E_{K}$, and $E_{L}$ are the reversal potentials for sodium, potassium, and leak channels.
The gating variables obey first-order kinetics and are described by the following differential equations:
$\frac{dm}{dt}$ $=\alpha_{m}(1-m)-\beta_{m}m$ (4)
$\frac{dn}{dt}$ $=\alpha_{n}(1-n)-\beta_{n}n$ (5)
$\frac{dh}{dt}$ $=\alpha_{h}(1-h)-\beta_{h}h$ (6)
Where $m$, $n$, and $h$ are the gating variables representing the activation and inactivation states of the ion channels. $\alpha_{m}$, $\alpha_{n}$, $\alpha_{h}$, $\beta_{m}$, $\beta_{n}$, and $\beta_{h}$ are voltage-dependent rate constants for the gating variables. These rate constants are defined by empirical equations based on experimental data. HH model accurately describes the electrophysiological processes of neuronal activity through mathematical equations, allowing scientists to understand the fundamental mechanisms of a group of neurons’ behavior at the molecular and cellular levels.
Another famous theory is cable theory[16]. Cable theory is a mathematical model used to understand how electrical signals propagate along neurons, especially in their dendrites and axons. This theory applies the principles of electrical circuits to describe how the passive properties of the neuron influence signal transmission. It’s particularly useful in explaining the conduction of graded potentials, such as synaptic potentials and electrotonic (passive) spread of membrane potential changes.
The core equation of cable theory is derived from the cable equation, which is a form of the second-order linear partial differential equation. It can be expressed as:
$\frac{\partial^{2}V}{\partial x^{2}}=\frac{1}{r_{m}}\left(\frac{\partial V}{\partial t}+\frac{V}{r_{c}}\right)$ (7)
, where $V$ is the membrane potential, $x$ is the distance along the cable, $t$ is time, $r_{m}$ is the membrane resistance per unit length, $r_{c}$ is the cytoplasmic (axial) resistance per unit length. This equation describes how the voltage $V$ changes along the length of the neuron $x$ and over time $t$.
Cable theory has several important applications in neuroscience, including understanding signal decay, neural modeling and synaptic integration. Despite its usefulness, cable theory has limitations. First, real neurons have complex geometries, which are simplified in cable theory. This can lead to inaccuracies in predicting real neuronal behavior. The complex structure of neuron should not be neglected. Besides, cable theory primarily addresses passive electrical properties and does not account for active properties like action potentials, which are crucial for neural communication.
Apart from that, researchers proposed many variants of neuron dynamic models[17][18]. These models serve to elucidate the physiological activity of neurons and contribute to the description of neuron behavior.
However, these models has some limitations: these equations often contain a multitude of parameters. In instances where measurement techniques fall short of quantifying each parameter individually, estimation becomes a necessity.
Spiking Neural Networks (SNNs)[19] is designed to optimize the parameters in neural dynamic models. Unlike their traditional counterparts, which process information through continuous values, SNNs utilize discrete events known as spikes. These spikes, brief in duration, mimic the communication method of neurons within the brain, facilitating the transmission of information.
One of the simplest and most commonly used models to describe neuron dynamics in SNNs[19] is the Leaky Integrate-and-Fire (LIF) model[20]. The LIF model captures the essential characteristics of spike generation in neurons. The membrane potential $V(t)$ of a neuron at time $t$ is described by the following differential equation.
$\tau_{m}\frac{dV(t)}{dt}=-[V(t)-V_{rest}]+RI(t)$ (8)
Where $\tau_{m}$ is the membrane time constant, which determines how fast the membrane potential decays towards the resting potential $V_{rest}$ in the absence of input. $V_{rest}$ is the resting membrane potential. $R$ is the membrane resistance. $I(t)$ is the input current at time $t$. When $V(t)$ reaches a certain threshold $V_{threshold}$, the neuron fires a spike, and $V(t)$ is reset to $V_{rest}$ or another predefined value. After firing, the neuron may enter a refractory period during which it cannot spike, regardless of the input.
SNNs have enabled researchers to simulate networks comprising millions of neurons, tackling a wide array of tasks ranging from image classification and natural language processing to controlling robotic arms[21][22][23][24].
Despite their advances, SNNs face significant challenges. One primary issue is that the SNN’s structure often draws inspiration from anatomically defined from functional brain networks rather than accurately replicating the intricate neural connections observed at the microscale. In terms of the connection method of artificial neurons, SNNs usually use a structure similar to that of ANNs. Consequently, these models resort to architectures akin to those found in artificial neural networks, hindering their ability to authentically emulate brain functions. Additionally, SNNs are marked by their low computational efficiency, with a single forward propagation frequently requiring several seconds to complete. These limitations underscore the difficulties in considering SNNs as the optimal approach for brain simulation in their current state.
As for the large scale brain simulation model aims to simulate the whole brain’s behavior, Eugene et al.[25] present the first large-scale brain simulation model based on experimental measures in several mammalian species. The model is founded upon the global thalamocortical anatomy derived from diffusion tensor imaging of the human brain. It incorporates multiple thalamic nuclei and a six-layered cortical microcircuitry, which is grounded in in vitro labeling and three-dimensional reconstruction of single neurons from the cat visual cortex. Additionally, the model simulates over 1,000,000 SNN neurons, calibrated to replicate known in vitro response patterns recorded in rats. It boasts nearly half a billion synapses, incorporating accurate receptor kinetics, short-term plasticity, and dendritic spike-timing-dependent synaptic plasticity (dendritic STDP). The model can demonstrate behavioral patterns that mimic normal brain activity. This model is a comprehensive attempt to simulate the whole neural system, which is a critical component of the brain involved in sensory processing and consciousness.
Eliasmith et al.[26] introduce a more comprehensive brain simulating model, which consists of 2,500,000 neurons and serves to close this divide by demonstrating a variety of behaviors. This model receives only sequences of visual images and produces all its responses through a simulated arm. Despite its simplifications of real structure, the model encapsulates numerous elements of neuroanatomy, neurophysiology, and behavioral psychology, as evidenced by its performance across eight distinct tasks.
As for the standardized platform, Zeng et al. [27] present BrainCog, a cognitive intelligence engine inspired by the brain. This platform, based on SNN, offers critical infrastructure for the creation of brain-inspired artificial intelligence and simulation of the brain. BrainCog encompasses a variety of biological neurons, encoding methods, learning principles, brain regions, and a combination of hardware and software as its core elements.
However, it is essential to note that each neuron’s dynamic model involves multiple differential equations, posing a significant challenge in simulating the complete behavior of the intact brain given the constraints of available computing power. Therefore, there is a compelling need for a cost-effective simulation approach to address this challenge.
As for the collective action of neuron populations, despite the success of neural system mean field models[28] in areas such as modeling epileptic seizures[29], there is still a lack of models that precisely describe the large-scale activity of whole brain.
2.2 Brain Network Communication Model
The network communication model is a conceptual framework that describes how data is transmitted from one node to another over a network. It breaks down the complex process of communication into simpler, more manageable layers or components, each with a specific role. The most widely recognized network communication models are the Open Systems Interconnection (OSI) model and the Internet Protocol Suite, commonly known as the TCP/IP model. Nowadays, scientists apply network communication models to neuron research, which are mostly used to understand how brain networks operate[30][31][32].
Early studies have indicated that, similar to social and electrical networks[33], brain networks also exhibit properties such as power-law distributions[34] and the rich-club phenomenon[35]. These results indicate that brain networks contain a wealth of network statistical characteristics. To reveal the mystery of the brain, researchers must take network characteristics into consideration.
Researchers designed many brain network models to describe the processes of message propagation in brain. One category of brain network communication models is the routing protocol model. The most famous model of this category is shortest path model[36], which assumes that signals between a pair of nodes travel along the shortest path between them. Let $d(v)$ represent the shortest path length from the starting node to node $v$, for each node $v$ in the network, initialize $d(v)=\infty$, then for the starting node $s$, set $d(s)=0$. Next, process each node in increasing order of $d(v)$, for each node, update the $d(u)$ value of its neighbors $u$. The update way can be represented as follows:
$d(u)=\min(d(u),d(v)+w(v,u))$ (9)
where $w(v,u)$ is the weight of edge $e_{uv}$.
However, the key disadvantage of the model is these methods require each neuron must have knowledge about the entire network structure[37], which doesn’t correspond to the reality: every neuron only has the message of its neighbor neurons.
[…]
However, the main disadvantage of navigation model is it supposes that neurons know the distance between their neighbors and a desired target region. Recently their still has a lack of evidences that reveal how this knowledge would be given to neurons in brain networks.
Another well-known category of models is the diffusion model. The diffusion model posits that signals propagate through random walk dynamics. This process of propagation does not require individual neurons to possess any prior knowledge beyond their immediate neighboring neurons. Among the most classic of these is the random walk model[40]. In the random walk model, signals are transmitted to randomly selected neighboring nodes with a probability proportional to the connection weights until they reach the target node. For unbiased random walk, the function of random walk model is as follows.
$P_{ij}=\frac{1}{k_{i}}$ (10)
Where $P_{ij}$ represent the probability of node $i$ choosing node $j$.$k_{i}$is the degree of node $i$.
As for biased random walk, for example, if the transition probability depends on the edge weights or certain properties of the nodes, then the transition probability can be adjusted to
$P_{ij}=\frac{w_{ij}}{\sum_{l\in N(i)}w_{il}}$ (11)
Where $w_{ij}$ is the weight of node $i$ to node $j$,$N(i)$is the set of node $i$.
However, studies have shown that the efficiency of signal transmission through random walks is relatively low[41]. So it can’t describe the rapid communication process in the brain.
An effective category is threshold model[42]. The threshold model was originally proposed to characterize collective behavior. In the threshold model, a node’s behavior depends on the behavior of its neighboring nodes. Suppose that within a group, nodes can hold one of two states regarding a certain behavior: an active state and a quiescent state. For a node, it will only transition from the quiescent state to the active state if the stimulus given by its neighboring nodes in the active state surpasses a certain threshold; otherwise, it remains in the quiescent state. The condition for node $i$ to adopt the behavior at time $t$+1 can be formulated as follows.
$a_{i}(t+1)=1\quad\text{if}\quad\frac{1}{k_{i}}\sum_{j\in N_{i}}a_{j}(t)\geq\theta_{i}$ (12)
Where $N_{i}$ is the set of neighbors of node $i$, $a_{i}(t)$ is the adoption state of node $i$ at time $t$. $\theta_{i}$ is the threshold of node $i$. $k_{i}$ is the degree of node $i$. The threshold model has been extensively applied in artificial neural networks.
However, how to set the propagation thresholds for different neurons, and how to design the dynamics model between neurons remains a challenge.
3 Data and Methods
3.1 Drosophila Connectome Data Set
We employ the recently released largest Drosophila brain connectome[14], offering a thorough mapping of the entire neural network within a Drosophila female brain. This data set encompasses over 120,000 neurons and 30,000,000 synapses, including detailed information on synaptic connections between neurons. Notably, it provides high-precision 3D coordinates, accurate to the nanometer level, for both cell bodies and synapses. Additionally, this data set offers extensive labeling for various types of neurons. For a more in-depth understanding, please refer to [14][43] for detailed information.
Table 1 shows the basic properties of the data set. Note that we combine multiple edges between two nodes to calculate network statistic properties. The average degree before combining multiple edges is 501.6 while average degree is 37.01 after combination. It means most of the pairs of neuron has many links.
Table 1: Statistical properties of Drosophila Brain Connectcome Graph
Number of Nodes Number of Edges Average Degree
$131459$ $32970606$ $37.01$
Average Degree (before combination) Clustering Coefficient Eigenvector Centrality
$501.6$ $0.1527$ $0.0008817$
Apart from that, we draw the curves of degree distribution. From Fig. 1 We found that the distribution of log(degree) initially appears to be approximately linear, but at the end it exhibits a long-tail distribution, which demonstrate that the original distribution of degree follows a power-law distribution. This result corresponds to the earlier research that neural network also has the power-law property[34].
Figure 1: Degree Distribution (Log-Log Scale)
Also, we plot the correlation curves between degree and cell surface area, volume, and cell length (See in Fig. 2). The results indicate the degree distribution is logarithmically linearly correlated with these data. These characteristics of the brain can explain Weber-Fechner’s law[44]: for any sensory pattern, the perceived intensity is a logarithmic function of the physical intensity. The structural changes can only cause a logarithmic change in the average connection strength.
Figure 2: Correlation between degree distribution and neuron characteristics
These statistic results are a valuable supplement to previous findings[45] and are crucial for a deeper understanding of this data set.
3.2 Model
In this section, we introduce the large scale network communication model which consumes less computational resource and can efficiently utilize the network structure. First we introduce the symbol systems in this article.
The original network presentation of neural systems often uses nodes to represent the neurons and using edges to represent the links (chemical synapses, electrical synapses, etc) between neurons[4]. However, this approach has a number of limitations: Firstly, there are more than one link (synapse) between two neurons[46] and the links are spatially and temporally asymmetric. To illustrate, two neurons shown in Fig. 3 have multiple synapses between them, and the dynamical properties of these different synapses may vary. Additionally, due to their spatial locations, the signal transmission delay and the position of action on the downstream neurons are also different, making it difficult to simply merge them.
Figure 3: An example of Complex Connection
For this reason, similar to previous research[47], we use the dual method on Drosophila connectome data set. The method is introduced below.
A dual graph is a concept in graph theory which describes a special relationship between graphs. In a dual graph, the nodes and edges of the original graph are interchanged, resulting in a new graph. In simple terms, a dual graph is obtained by replacing the nodes of the original graph with edges, and the edges with nodes. Fig. 4 is an example of an original graph and a dual graph. After being processed, the new graph can be considered to be a hyper graph, whose nodes represent the synapses and links represent the cell body. Then we can use the hyper communication model to describe it.
(a) Original Graph
(b) Dual Graph
Figure 4: An example of original graph and dual graph
In order to standardize the symbol system, we use a set to represent the neuronal cell body and use an element to represent a physical link between neurons. We regard the process of messages interactions as the elements’ interactions on sets.
Let $N=\{n_{1},...,n_{i},...\}$ represent the set of all elements and $E=\{e_{1},...,e_{j}...\}$ represent the probable sets of elements. Note that the intersection of sets include some elements, which is formalized as the following manner.
$e_{k}\cap e_{q}=\{n_{i}|(n_{i}\in e_{k})\wedge(n_{i}\in e_{q})\}$ (13)
Note that union of every element in $N$ and $E$ is the whole graph. Let $G$ represent the graph. The graph’s definition is given as follows.
$G=\bigcup_{j}e_{j}=\bigcup_{i}\{n_{i}\}.$ (14)
Because neural link is directed, let ${e_{j}}^{+}$ ( ${e_{j}}^{-}$) represent the set of elements which output (input) the signals, note that ${e_{j}}^{+}$ = ${e_{j}}^{+}\cup{e_{j}}^{-}$.
Every neuron and synapse has a activate state (state = 1) or resting state (state = 0). So set the statement of the element $i$ as $X_{i}$. $X_{i}=1$ if the element $n_{i}$ is active, otherwise $X_{i}=0$. Then let $Y_{j}$ represent the statement of set $e_{j}$. $Y_{j}=1$ if the set $e_{j}$ is active, otherwise $Y_{j}=0$.
We use a linear threshold activation function. Then, general transmission function is defined in the following manner,
$Y_{j}=\sum_{\{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset\}}\sum_{n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\sigma(\omega(n_{k},e_{i},e_{j})X_{k})$ (15)
where $\omega(n_{k},e_{i},e_{j})$ represents the link weight between set $i$ and set $j$ (specifically, two neurons). Here we consider that the link weight between two neurons is not only related to the neurons themselves, but also to the synapses between them. $\sigma(.)$ is the activate function.
The statement of an neuron will influence the statement of all output synapses. The equation between neuron $j$ and output synapse $i$ is as the following manner.
$X_{i}=Y_{j}$ (16)
The link weights between neurons determine the extent of influence they have on each other. Larger link weights mean a tighter ”link” between two neurons. In everyday life, the brain continuously changes the link weights between neurons in response to inputs from the external world, allowing intelligence to emerge from the micro to the macro. Therefore, if researchers want to simulate the process of information propagation in the brain on large scale in a computer, determining the link weights between neurons is essential. However, existing experimental methods are not well-suited for directly measuring the weights between neurons in a whole brain, so adopting some alternative approach is necessary. Below, we introduce two hypotheses that use different methods to estimate the connection weights between neurons.
3.2.1 Average Hypothesis
As previously mentioned, there exists a large number of synapses between two neurons. However, in practical applications, it is difficult to set the parameters for these edges individually. Consequently, a simple assumption naturally arises: assuming that the source neuron has the same influence on the target neuron through every synapse. We propose an assumption that every synapse is equivalent, which is widely used in neuron graph statistics analysis[45]. That is to say,
$\omega(n_{k},e_{i},e_{j})=1$ (17)
Then the normalized equation is as follows.
$Y_{j}=\sigma(\frac{\sum_{\{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset\}}\sum_{n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\sigma(\omega(n_{k},e_{i},e_{j})X_{k})}{{{Max}(\{\omega(n_{k},e_{i},e_{j})|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset,n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\})})$ (18)
3.2.2 Distance Hypothesis
The continuous evolution of organisms has led to the brain’s spatial structure being highly complex, with a small area possibly containing a large number of neurons. Besides, a fact is messages are transmitted through bioelectric currents. Assuming that the distance is longer, the resistance will be greater, resulting in a smaller communication effect. Based on this, there emerges a hypothesis: Is the strength of interactions between neurons inversely proportional to the spatial distance between them? So in the distance hypothesis we assume that the impact of input synapses on neurons is inversely proportional to the distance between them. Then the norm can be formulated as follows.
$\omega(n_{k},e_{i},e_{j})=(\parallel\vec{r_{k}}\parallel^{2}-\parallel\vec{r_{j}}\parallel^{2})^{\frac{1}{2}}$ (19)
Where ${r_{k}}$, ${r_{j}}$ are the center position vector of synapse and cell body, respectively.
Then the transmission function is defined as the following manner.
$Y_{j}=\sigma(\frac{\sum_{\{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset\}}\sum_{n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\sigma(\omega(n_{k},e_{i},e_{j})X_{k})}{{{Max}(\{\omega(n_{k},e_{i},e_{j})|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset,n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\})})$ (20)
3.3 Evaluation Criteria
In previous study, researchers used classic indicators such as signaling cost and activation time to evaluate the quality of the model[11][48][49]. However, these indicators can’t prove the model accords with the true communication process even if the activation time is short and the signaling cost is minimum, etc. To better assess whether a model can reflect the real situation of brain communication, we use the percentage of activated neurons in regions that should work under a certain stimulus. To illustrate, if the brain gets a visual stimulus, the region related to vision should work while the region not related to vision should keep in a resting state. Then we record the percentage of neurons activated and if regions related to the certain function have a significant higher activation percentage than the average activation percentage of the whole brain, we call it a quasi-real activation pattern.
We define the consistency indicator $C$. The consistency indicator $C$ in region $r$ which has a certain function (visual, olfactory, etc) is defined in the following manner.
$C_{r}=N_{r}^{act}/N_{r}$ (21)
Where $N_{r}^{act}$ represents the number of neurons activated in area $r$ at the final state, $N_{r}$ represent the number of neurons in area $r$ and $N$ represent the total number of neurons.
3.4 Large Spatial Network Visualization
The visualization software of experiment observation on large spatial networks is important for researchers[50][51]. However, there is still a lack of visualization system that can observe information propagation on large spatial networks in real-time, which poses difficulties in experiment. To observe directly the communication progress on this super large Drosophila connectome, we develop a large spatial network visualization software.
The drawing part mainly employs the Canvas provided by HTML5 and we use the Three.js based on WebGL technology to achieve spatial network visualization effects, while the user interface display is constructed by Vue.js, providing users with an intuitive and interactive interface to display and analyze data. Fig. 5 is an example of the 3D visualization system. This software will be made available to other researchers with the author’s permission.
(a) Side view 1
(b) Side view 2
(c) Side view 3
(d) Top view 1
(e) Top view 2
(f) Top view 3
(g) Front view 1
(h) Front view 2
(i) Front view 3
Figure 5: Visualization of Spatial Drosophila Brain Network
3.4.1 Front End Technical Details
We use the Vue.js framework to build the user interface. Vue.js is a popular JavaScript framework that simplifies the development of web applications through componentization, enabling the creation of front-end interfaces with rich interactivity and good user experience.
The rendering of three-dimensional complex networks relies heavily on the Three.js library. Three.js is a powerful JavaScript library for creating and displaying 3D graphics. It offers a rich set of 3D objects, materials, lights, and animations, allowing for the easy creation of high-quality 3D scenes and objects.
To improve rendering performance, GPU acceleration technology is utilized. GPU acceleration takes advantage of the computational power of the Graphics Processing, significantly enhancing rendering speed and efficiency through parallel processing and stream processing modes. The GPU acceleration capabilities of Three.js, especially through the creation of point objects using Float32BufferAttribute, which contains their position information in three-dimensional space, greatly enhance the rendering speed and smoothness.
3.4.2 Back End Technical Details
We use the Spring Boot framework, an open-source framework based on Java for quickly building independent, production-grade Spring applications. With Spring Boot, backend services can be rapidly set up to provide RESTful API interfaces for frontend consumption. The backend also utilizes the PostgreSQL database to store network data.
Besides, We use PostgreSQL, which is a powerful open-source relational database management system that offers a rich set of data types and robust querying capabilities. To store 3D network data, we leverage PostgreSQL’s JSONB type, which allows storing JSON-formatted data within PostgreSQL and provides a series of functions and operators for querying and manipulating this data. With the JSONB type, information about nodes and edges in the network can be stored, enabling fast querying and updating operations.
4 Experimental Results and Analysis
In this part we introduce the experimental results and analysis. Above all, we introduce the activation pattern experiment. We give different stimuli and record the activation pattern of different neurons. In addition, we give a unilateral stimulus (stimulus one hemisphere) and record the bilateral response of neurons. Finally, to prove the network structure matters in brain network communication, we calculate the average physical and network distance between different areas.
4.1 Experimental Settings
The basic methodology of the experiment is described as follows. In this study, a continuous stimulus signal (binary 0-1 signal) is applied to a subset of input neurons (such as visual or olfactory neurons) in the brain. Subsequently, we compute the signal’s propagation process within the brain and record the responses of both intermediate and output neurons. The evaluation criterion is determined by calculating the ratio of neurons that ought to be activated to those that are actually activated in a specific area. Given that the brain functions as a type of shallow neural network[52], we document the outcomes at an iteration step size of 5, by which point the majority of neurons have achieved a state of stability. Moreover, as the input remains constant and no randomization parameters have been introduced, the outcomes of each experiment are consistent, eliminating the need for repetition.
Next, we will introduce the background knowledge of the neurons involved in this paper. Here are introductions of these neurons.
Optic neurons, also known as retinal ganglion cells, are the first neurons in the visual pathway. They receive signals from the photoreceptor cells in the retina and transmit this information via the optic nerve to the brain. These neurons play a crucial role in carrying visual stimuli from the eye to the brain, where further processing and interpretation of the visual information occur.
Visual projection neurons are a type of neuron found in the visual system that transmit visual information from one region of the brain to another. They play a crucial role in processing and relaying visual signals from the retina to various visual centers in the brain, including the thalamus, hypothalamus, and primary visual cortex (V1), among others.
Antennal Lobe Input Neurons (ALINs) are primarily olfactory sensory neurons (OSNs) that carry olfactory information from the sensory structures, such as the antennae and maxillary palps, directly to the antennal lobe. Each OSN expresses a specific odorant receptor and responds to particular chemical cues. The axons of these neurons synapse with projection neurons and local neurons within the glomeruli of the antennal lobe.
Antennal Lobe Projection Neurons (ALPNs) transmit the olfactory information in the antennal lobe. These neurons receive input from the olfactory sensory neurons within the glomeruli and project their axons to other brain areas, such as the mushroom body and the lateral horn, where further processing and integration of olfactory information occur.
Antennal Lobe Output Neurons (ALONs) refer to the projection neurons that serve as the primary output pathways of the antennal lobe, sending processed olfactory information to higher brain regions.
Antennal Lobe Local Neurons (ALLNs) are inter neurons that are confined to the antennal lobe. They typically have processes (dendrites and axons) that branch extensively within the antennal lobe and form connections with multiple glomeruli. These neurons play a critical role in modulating and refining the olfactory information by providing inhibitory or excitatory input to projection neurons and other local neurons, thus shaping the olfactory responses and contributing to odor discrimination and perception.
For further information, please read [43].
4.2 Activation Patterns on Different Type of Neurons
Primarily, to validate the efficacy of our model in modifying signal transmission between input and output neurons, we showcase the activation patterns of intermediate and output neurons throughout the information transmission process. This demonstration serves to substantiate that our model accurately reflects real-world scenarios, shaping a quasi-real activation pattern.
The specific idea for the experiment is following : we stimulate different types of neurons (such as visual input neurons, olfactory neurons, etc.), and record the rate that neurons that are activated in different brain areas. If a specific stimulus causes a significantly greater response than the average in the brain area that it is theoretically supposed to activate, then it is considered that the model can generate a quasi-real activation pattern in the brain. We give 2 types of input neurons (visual,olfactory) stimulus display 6 type of neurons: optic (visual), visual_projection (visual), ALLN (olfactory), ALIN (olfactory), ALPN (olfactory) and ALON (olfactory).
To demonstrate the importance of network structure, we add a comparative model that randomly reconnects networks. Specifically, we calculate the average network distance from input neurons (visual/olfactory) to corresponding neurons (optic, visual_projection/ALLN, ALPN, ALON, ALLN), round it to the nearest integer to determine the average layer of that neuron, and then randomly swap the connections between nodes within each layer.
Fig. 6 shows the activation patterns of different models. The different neuron types are on the horizontal ordinate and the activation percentage is on the vertical ordinate. The horizontal line represents the percentage of activated neurons in the whole brain. From Fig. 6 we can find that under the average hypothesis, visual stimulus cause a strong impact (significantly more than average) on visual neurons (optic, visual projection) while causing a slight impact (significantly less than average) on other type of neurons. Specifically, under visual stimulus, in response to visual stimuli, the optic (vision related area) exhibit a 10.26% higher activation rate compared to the average level, while no olfactory-related neurons are activated. In addition, olfactory stimulus cause a strong impact (significantly bigger than average level) on olfactory neurons (ALLN, ALIN, ALPN, ALON), exceeding the average level by 81.49%. And it cause a slight impact (significantly less than average) on other type of neurons. For example, it only activates 0.04% of optic and 0.79% of visual_projection.
Furthermore, the results are not as significant on the distance model. Some incorrect areas were activated: The activate rate of ALLN is only 7.08%, with a 5.16% difference from the average level (Fig. 6(c)). Besides, ALIN, ALPN and ALLN, which should not be activated, have activate rates of 12.50%, 0.15%, and 1.89%, respectively (Fig. 6(c)). These results indicate that whether distance hypothesis can describe the actual situation remains to be discussed, which is a refutation for previous research[53].
As for the shuffle model, the ALIN, ALPN, ALON and ALLN which should not be activated have activate rates of 70.00%, 54.13%, 61.25% and 62.31%, respectively (Fig. 6(e)). Although these activate rates do not meet the average level, they are significant activation rates that means many neruons are activated incorrectly. Moreover, optic and visual_projection are activated incorrectly, with activate rates of 55.11% and 42.52%, respectively (Fig. 6(f)). Besides, ALLN, ALPN, ALON, and ALLN which are supposed to exceed the average level are actually lower than the average level by 0.01%, 17.74%, 21.25% and 9.34%, respectively (Fig. 6(f)).
These results indicate that after shuffling the network structure, the original quasi-real activation patterns completely disappear. It means the real network structure matters in generating the quasi-real activation patterns. In addition, even just designing simple models, it is feasible to simulate the propagation of information in the brain at the level of neurons and generate a quasi-real activation pattern.
(a) visual stimulus, average model
(b) olfactory stimulus, average model
(c) visual stimulus, distance model
(d) olfactory stimulus, distance model
(e) visual stimulus, shuffle model
(f) olfactory stimulus, shuffle model
Figure 6: Experiment of intrinsic neurons’ activation. $\sigma_{0}=0.8$. For shuffle model, we conduct 5 repeated experiments and take the average and standard deviation.
4.3 Distances between Input and Activation Areas
In the previous section, the quasi-real activation patterns on different type of neurons of the brain was presented. Through observation, this relationship was identified, and a quasi-real activation pattern was observed in the drosophila brain. There is an issue here: Is the quasi-real activation pattern determined by the brain network’s structure, the spatial structure of the brain, or the quality of the network communication model? Based on the puzzle, we calculated the average network distance and spatial distance between different areas of the brain network. The optic and visual_projection area are related to vision and ALLN, ALIN, ALON, ALPN, LHCENT and LHLN are related to olfactory. The results are presented in Tab. 2.
From Tab. 2, it is evident that neurons which are closer in spatial distance are not necessarily easier to activate. For example, all areas and visual input neurons have similar average spatial distances, but the quasi-real activation patterns can only be observed in specific neural types (optic, visual_projection). In contrast, the average network distance between input neurons and neurons should be activated is closer than irrelevant neurons. For instance, the distance between visual input neurons and optic, visual_projection are 5.7018 and 5.7432, respectively. And the distance between visual input neurons and ALLN, ALIN, ALON, ALPN, LHCENT, LHLN which are irrelevant with vision are 6.6442, 6.2350, 6.1506, 6.5380, 6.3094 and 6.8206, respectively. The network distance of neurons related to visual input neurons is generally closer than the distances of unrelated neurons. Besides, The average network distance of neurons related to olfactory input neurons is approximately twice as close as the distance of unrelated neurons. These results indicates that compared to spatial distance, the average network distance between input neurons and their corresponding activation areas is closer, meaning that compared to spatial structure, network structure is more important, determining these quasi-real activation patterns of the brain network.
This results reflect the importance of network structure in neural networks, that is, the ability to reproduce certain functions of the brain by considering only the network structure.
Table 2: Spatial and network distance between input and main areas. The unit of spatial distance is nm and the unit of network distance is the number of edges in the shortest path.
distance type Neuron type Area Name optic visual_projection ALLN ALIN
Spatial Distance Visual Input Neurons $365012$ $354725$ $351919$ $348901$
Olfactory Input Neurons $302690$ $251144$ $110827$ $173345$
ALON ALPN LHCENT LHLN
Visual Input Neurons $346925$ $348517$ $362825$ $366013$
Olfactory Input Neurons $123179$ $115089$ $212515$ $246719$
optic visual_projection ALLN ALIN
Network Distance Visual Input Neurons $5.7018$ $5.7432$ $6.6442$ $6.2350$
Olfactory Input Neurons $5.7967$ $5.1706$ $3.4458$ $3.3845$
ALON ALPN LHCENT LHLN
Visual Input Neurons $6.1506$ $6.5380$ $6.3094$ $6.8206$
Olfactory Input Neurons $3.6312$ $3.6820$ $3.9557$ $3.9758$
4.4 Bilateral Response under Unilateral Stimulus
In real life, the left and right hemispheres cooperate to perform complex tasks. We are interested in whether network models can simulate this cooperative effect to some extent. So we administer a unilateral stimulus to the input neurons and monitor the varying activation ratios across both hemispheres. More precisely, we independently simulate visual input neurons in the left and right hemispheres and record the rate of activated neurons in each hemisphere over time. Figure 7 illustrates the curve depicting how the response ratio evolves with each time step. It’s noteworthy that when simulating one hemisphere alone, visual processing neurons in both hemispheres are activated. At time step 5, the unstimulated hemisphere begins to have a activate pattern, which corresponds to the hypothesis of shallow connection[52]. Furthermore, our observations reveal a concurrent effect: the activation pattern in one hemisphere mirrors that in the unstimulated hemisphere.
These findings are reflective of reality. Although the left and right hemispheres of the brain possess certain functional specializations, they also demonstrate a synergistic effect. Typically, they collaborate to fulfill specific functions[54]. For instance, during complex cognitive tasks, both hemispheres coordinate their efforts to contribute to task completion. Even when simulating input to neurons on just one side of the brain, our model reveals similar synergistic effects, indirectly attesting to the model’s effectiveness. Besides, the mirror effect align with real situation, underscoring the brain’s resilience as a system. In instances of disease or physical damage leading to impairment in one hemisphere, this interconnected structure supports the continuation of critical functions, such as vision.
(a) Left stimulus
(b) Right stimulus
Figure 7: Experiment of activation rate under single stimulus. We set threshold $\sigma_{0}=0.8$.
Visually, the shape of the brain appears to be physically similar, but whether the network structures are similar remains unknown. To explore the principle behind similar responses caused by a unilateral stimulus, we calculated the degree, clustering coefficient, and eigenvector centrality distribution for each hemisphere. The Pearson correlation coefficient of degree, clustering coefficient, and eigenvector centrality vector is 0.9986, 0.9988, and 0.9991, respectively. According to Fig. 8, these statistical indicators for both hemispheres are extremely similar, indicating that the brain exhibits strong symmetry in network structure.
(a) Degree
(b) Clustering
(c) Eigenvector centrality
Figure 8: Degree, clustering coefficient, and eigenvector centrality distribution
5 Conclusion
Considering the insufficient evidence supporting the vitalness of network structure in brain message transmission and the effectiveness of network models in simulating brain activation patterns, we propose a large scale network communication model based on simple propagation rules. We also design an evaluation criteria to compare real and simulated activation patterns. Our research utilizes the largest adult Drosophila connectome data set, where we analyze the basic statistical properties and explore the relationship between neuron length, area, size, and connectivity degree. Experimental results demonstrate that our model produces an activation pattern across the brain that closely resembles reality when subjected to identical stimulus on input neurons. This result strongly suggests that even with simple propagation rules, network models can approximate the brain’s activation patterns. Besides, when changing the network structure, there is an incorrect activation pattern on the brain network. It is an evidence that network structure matters in generation real brain activation patterns. Furthermore, to prove network structure matters rather than propagation rules or spatial structure, we assess the network and spatial distances across different areas, discovering that the average network distance between input neurons and their corresponding activation areas is shorter. Notably, applying unilateral stimulus to input neurons elicit bilateral responses, mirroring bilateral brain coordination in reality and providing additional evidence that simple network models can closely replicate reality. Moreover, we calculate the statistic properties and find that the network structures of two hemispheres are extremely similar. We also make a large spatial network visualization software. This research reveals network models can reach the brain’s quasi-activation pattern even with simple propagation rules. Besides, it provides evidence that network structure matters in brain activity pattern generation. Other artificial neural network research may need to make their own structure close to the real network structure. However, it does not perfect the model to achieve a full-activation pattern. Future research could enable complete simulation of brain behavior by optimizing propagation rules and automatically adjusting link weights, paving the way for achieving genuine artificial intelligence.
6 Acknowledgement
This work was supported by the National Natural Science Foundation of China (72025405, 72088101), the National Social Science Foundation of China (22ZDA102), the Hunan Science and Technology Plan Project (2020TP1013, 2020JJ4673, 2023JJ40685), the Innovation Team Project of Colleges in Guangdong Province (2020KCXTD040), and the Postgraduate Scientific Research Innovation Project of Hunan Province (XJCX2023095). The authors declare that they have no conflict of interest.

### v3

@@
-Simple Network Mechanism Leads to Quasi-Real Brain Activation Patterns with Drosophila Connectome
+Network Structure Trumps Neuron Dynamics: Insights from Drosophila Connectome Simulations
 Abstract
-Considering the high computational demands of most methods, using network communication models to simulate the brain is a more economical way.
-However, despite numerous brain network communication models, there is still insufficient evidence that they can effectively replicate the real activation patterns of the brain.
-Moreover, it remains unclear whether actual network structures are crucial in simulating intelligence.
-Addressing these issues, we propose a large scale network communication model based on simple rules and design criteria to assess the differences between network models and real situations.
-We conduct research on the biggest adult Drosophila connectome data set.
-Experimental results show significant activation in neurons that should respond to stimulus and slight activation in irrelevant ones, which we call quasi-real activation pattern.
-Besides, when we change the network structure, the quasi-activation patterns disappear.
-Interestingly, activation regions have shorter network distances to their input neurons, implying that the network structure (not spatial distance) is the core to form brain functionality.
-In addition, giving the input neurons a unilateral stimulus, we observe a bilateral response, which is consistent with reality.
-Then we find that both hemispheres have extremely similar statistical indicators.
-We also develop real-time 3D large spatial network visualization software to observe and document experimental phenomena, filling the software gap.
-This research reveals network models power: it can reach the quasi-activation pattern even with simple propagation rules.
-Besides, it provides evidence that network structure matters in brain activity pattern generation.
-Future research could fully simulate brain behavior through network models, paving the way for artificial intelligence by developing new propagation rules and optimizing link weights.
+Despite the success of artificial neural networks, the necessity of real network structures in simulating intelligence remains unclear.
+Utilizing the largest adult Drosophila connectome data set, we constructed a large-scale network communication model framework based on simple neuronal activation mechanisms to simulate the activation behavior observed in the connectome.
+The results demonstrate that even with simple propagation rules, models based on real neural network structures can generate activation patterns similar to those in the actual brain.
+Importantly, we found that using different neuronal dynamics models, all produced similar activation patterns.
+This consistency across different models emphasizes the crucial role of network topology in neural information processing, challenging views that rely solely on neuron count or complex individual neuron dynamics.
+Moreover, we test the influence of network reconnect rate and find that at least 1%’s reconnect rate will ruin the activation patterns appeared before.
+By comparing network distances and spatial distances, we found that network distance better explains the information propagation patterns between neurons, highlighting the importance of topological structure in neural information processing.
+To facilitate these studies, we developed real-time 3D large spatial network visualization software, bridging a crucial gap in existing tools.
+Our findings underscore the importance of network structure in neural activation and provide new insights into the fundamental principles governing brain functionality.
 Keywords: brain network, network communication model, drosophila connectome, activation pattern
 1 Introduction
-The neural systems of intelligent beings maintain constant communication[1][2].
-The intricate mechanism of this communication unfolds as follows: specific neurons, often input neurons, receive stimulus signals from the external environment or other parts of the body, transitioning from a resting to an activated state.
+The neural systems of intelligent beings maintain constant communication[1, 2].
+The intricate mechanism of this communication unfolds as follows: specific neurons, often input neurons, receive stimulus signals from the external environment or other parts of the body, changing from a resting to an activated state.
 Subsequently, they relay these stimulus signals to the next tier of neurons through synapses.
 This process gives rise to macroscopic emergence across the entire neural system.
-Such communication fosters the emergence of intelligence in all intelligent life forms, including humans.
+Such communication fosters the emergence of intelligence in all intelligent life forms, including human beings.
 Consequently, comprehending and replicating this process has become an indispensable avenue for unraveling biological intelligence and realizing artificial intelligence[3].
 Neural systems are interconnected networks[4], with neurons engaging in constant communication via synapses within the network’s topology.
-The brain’s unique and complex network structure is pivotal to the emergence of intelligence[5].
-Many scientists posit that network science is crucial for deciphering the mechanisms underlying the flexible regulation of neural communication[6][7][8].
+Some scientists think the brain’s unique and complex network structure is pivotal to the emergence of intelligence[5].
+Many scientists posit that network science is crucial for deciphering the mechanisms underlying the flexible regulation of neural communication[6, 7, 8].
 Determining how to leverage network science methodologies stands as a paramount challenge in contemporary neuroscience.
-Within this discipline, neural networks are often represented as graphs, consisting of nodes (entities) and edges (relationships).
-From the broadcasting model[9] to the diffusion model[10], researchers strive to craft network models that mimic activation patterns observed in the brain, offering insights into the workings of neural networks.
-However, despite artificial neural networks (ANNs)’ deviation from the brain’s actual structure, they get full achievements in achieving intelligence, which cast doubt on the necessity of real network structures in brain behavior simulation.
-Furthermore, some researchers have highlighted the absence of evidence linking network model simulations to genuine brain activity patterns[11], undermining the models’ status as a robust research tool in neuroscience.
-The effectiveness of network models in replicating the neuronal system’s activation dynamics remains uncertain.
-This uncertainty partly stems from many models’ failure to replicate the brain’s structure at the neuron and synapse level, compromising the models’ interpretability.
-Moreover, current brain network communication models often rely on prevalent metrics such as signaling cost and activation time to assess model quality[11].
+However, whether this network structure determines the emergence of intelligence remains a question.
+In recent years, artificial neural networks (ANNs) have taken rapid development in enormous fields[9, 10, 11].
+A notable recent achievement is large language model (LLM).
+However, the structure of any ANNs does not follow the actual brain network structure strictly, which poses a challenge to the viewpoint that network structure inherently contains intelligence.
+Simultaneously, many scientists in machine learning adhere to the scaling law: the more parameters the network has, the greater the degree of emergent intelligence[12].
+This viewpoint suggests that intelligence is solely related to the number of parameters a model has, rather than the unique network structure between neurons.
+Under this perspective, the reason biological brains generate intelligence is due to the number of neurons, and the unique network structure is a coincidence of evolution.
+To demonstrate the importance of brain network structure, network scientists are constantly exploring and experimenting.
+An approach is brain network communication model[13].
+Scientists wants to use brain network communication model to understand and describe the interaction and communication between different parts of the brain.
+This model aids in studying the structural and functional aspects of neuronal networks and how they coordinate to accomplish complex cognitive processes[14].
+The first challenge of this approach is obtaining a complete and accurate mapping of the brain’s structure.
+For example, current imaging technologies, such as MRI, fMRI, and electron microscopy, have limitations in resolution and speed.
+While MRI and fMRI provide non-invasive imaging, their spatial resolution is not sufficient to map individual neurons and synapses.
+On the other hand, techniques like electron microscopy, which can achieve higher resolution, are extremely slow and generate enormous amounts of data that are challenging to process[15].
+Modeling from the micro scale is necessary because modeling at the level of brain regions or nerve fibers does not include the most basic structural information and lacks the most complete physical foundation.
+Moreover, current models often rely on prevalent metrics such as signaling cost and activation time to assess model quality[16].
 Yet, these metrics fall short of verifying whether a network model accurately reflects the brain’s actual activation patterns, which is the fundamental goal of brain simulation.
 There is a need for evaluation criteria that can affirm whether network communication models genuinely mirror the brain’s communication processes.
-Additionally, observing the brain network communication process is important for researches.
-The absence of 3D visualization software capable of real-time monitoring of communication status and displaying networks with a vast number of nodes poses challenges for researchers, including neuroscientists, by impeding real-time observation, display, and analysis of experimental outcomes.
-Over the past few decades, biologists have dedicated themselves to determining the precise structure of the connectome, ranging from C.
-elegans[12] to Drosophila larvae[13].
-By understanding the connections within an organism’s nervous system, researchers can gain insights into how the nervous system operates.
-A recent study[14] has published the largest adult Drosophila connectome, providing data support for researches in network science.
-For these reasons, we conducted experiments utilizing the recently released comprehensive data set of the largest adult Drosophila connectome, enhancing the interpretability of our findings.
-Our approach involves crafting a large scale network communication model based on simple mechanisms, designed to simulate the activation behavior observed in the connectome.
+Based on facts and challenges above, benefiting from a largest-to-date adult Drosophila connectome data[17], this study attempts to investigate how the type of structure of neurons could affect the activation patterns of a brain.
+Our approach involves crafting a large scale network communication model framework based on simple mechanisms, designed to simulate the activation behavior observed in the connectome.
 Specifically, we establish the network representation of the connectome.
-Then we design each neuron’s state and threshold functions.
-When a neuron’s state meets the threshold, it activates and subsequently transmits signals to other neurons.
-After a few iteration, the whole network emerges an activation pattern.
+Then we design a large scale threshold propagation model of nerve cells: when a neuron received enough stimuli, it activates and subsequently transmits signals to other neurons.
+After a few iterations, the whole network emerges an activation pattern.
+We use threshold model and sigmoid model as the basic neuron model.
+We also design an improved LIF model as the basic neuron model for comparison.
 Additionally, we propose an evaluation criteria based on the ratio of actually activated neurons to the expected number of activation.
-Experiments reveals that our model has significant activation in the type of neurons that should be affected by the stimulus and slight activation in irrelevant neurons which called a quasi-real activation pattern.
-To prove the vitalness of network structure, we calculate every node’s layer and shuffle part of the network structure in same layer and find the shuffled network can’t generate any quasi-real activation pattern.
-To enhance our understanding, we developed a 3D network visualization software, utilizing HTML5’s Canvas and Three.js based on WebGL technology for large spatial network visualization effects.
+Experiments reveals that, from threshold , sigmoid, and LIF neuron model, no matter activation mechanism is, our model has significant activation in the type of neurons that should be affected by the stimulus and slight activation in irrelevant neurons.
+Moreover, to supply more evidence of the vitalness of network structure, we disturb the network by the way of degree-preserving rewiring and and find the disturbed network can’t generate any correct activation pattern.
+In addition, we test the sensitivity of disturb degree.
+For details, we record the activation rate of areas under different reconnect rates and find that 1%’s random change of network structure will cause a complete change in function.
+To prove the network structure, rather than the spatial structure, matters in activation pattern generation, we calculate the average network and spatial distance of input and activation neurons and results reveal that the average network distance between input neurons and their corresponding activation areas is shorter than unrelated areas, which is a strong evidence of network matters in message propagation in brain.
+Furthermore, when providing unilateral stimuli to input neurons, we observe a bilateral response in the Drosophila connectome, aligning with reality.
+Then we calculate the statistics properties and find that the left and right hemispheres have extremely similar degree, clustering coefficient, and eigenvector centrality vector coefficient.
+The vectors’ Pearson coefficient are up to 0.9986, 0.9988, and 0.9991, respectively.
+We also developed a 3D network visualization software see in appendix F, utilizing HTML5’s Canvas and Three.js based on WebGL technology for large spatial network visualization effects.
 This tool enables real-time observation of each neuron’s communication process.
-To prove the network structure matters (rather than spatial structure) in activation pattern generation, we calculate the average network and spatial distance of input and activation neurons and results reveal that the average network distance between input neurons and their corresponding activation areas is shorter than unrelated areas, which is a strong evidence to prove network structure matters in the process of message transmission in brain.
-Furthermore, by subjecting input neurons to the same stimulus, we observe significant responses in neurons expected to be affected, with only slight responses in irrelevant neurons—characterized as a quasi-real activation pattern.
-Notably, when providing unilateral stimuli to input neurons, we observe a bilateral response in the Drosophila neuronal wiring diagram, aligning with reality.
-Then we calculate the statistics properties and find that the left and right hemispheres have extremely similar degree, clustering coefficient, and eigenvector centrality vector coefficient.
-The vectors’ Person coefficient are up to 0.9986, 0.9988, and 0.9991, respectively.
-2 Related Works
+2 Results
+First, we report the basic statistics of Drosophila connectome.
+After that, to test the performance of our model in terms of realism on the data set and determine whether different artificial mechanisms can achieve similar activation patterns when the network structure changes, we give different stimuli and record the activation pattern of different neurons.
+In addition, we give a unilateral stimulus (stimulus one hemisphere) and record the bilateral response of neurons.
+Besides, to add more neurons’ real dynamic details, we also add an improved LIF model as the basic neuron dynamic model and record the activation patterns under stimuli.
+Finally, to prove the network structure matters in brain network communication, we calculate the average physical and network distance between different areas.
+2.1 Experimental Settings
+The basic methodology of the experiment is described as follows.
+In this study, a continuous stimulus signal (binary 0-1 signal) is applied to a subset of input neurons (such as visual or olfactory neurons) in the brain.
+Subsequently, we compute the signal’s propagation process within the brain and record the responses of both intermediate and output neurons.
+The evaluation criterion is determined by calculating the ratio of neurons that ought to be activated to those that are actually activated in a specific area (see as Equ.26 in appendix B).
+Given that the brain functions as a type of shallow neural network[53], we document the outcomes at an iteration step size of 5.
+Moreover, as the input remains constant and no randomization parameters have been introduced, the outcomes of each experiment are consistent, eliminating the need for repetition.
+For messages in related neurons, please read appendix C.
+In linear threshold model, we set $\sigma_{0}=0.8$.
+In the simulation of the LIF model, numerous hyperparameters need to be used.
+All the hyperparameters used are listed in the following table.
+Biological hyperparameters are set according to real biologic experiment test[54].
+Current proportion const is the optimal parameter obtained under equidistant search.
+Table 1: List of Hyperparameters
+Hyperparameter Description Value
+$\tau$ Membrane time constant 15.0 ms
+$V_{\text{rest}}$ Resting membrane potential -65 mV
+$V_{\text{th}}$ Spike threshold -45 mV
+[…]
+$dt$ time step 0.55 ms
+$\alpha$ Current proportion const 0.03/0.045 (visual/olfactory)
+2.2 Visualization and Statistic Properties of Drosophila Connectome
+In this section we first present the visualization of Drosophila Connectome (Fig.1(a)).
+Other visualization cases and methods of visualization are presented in appendix F.
+Visually, the shape of the brain appears to be physically similar, but whether the network structures are similar remains unknown.
+So we further investigate the network statistics of two hemispheres.
+We calculated the degree, clustering coefficient, and eigenvector centrality distribution for every node in each hemisphere (Fig.1(d), (e) and (f)).
+The Pearson correlation coefficient of degree, clustering coefficient, and eigenvector centrality vector in each hemisphere is 0.9986, 0.9988, and 0.9991, respectively.
+According to Fig.1(d), (e) and (f), these statistical indicators for both hemispheres are extremely similar, indicating that the brain exhibits strong symmetry in network structure.
+From Fig.1(c) we can see that the degree, i.e., the number of connections to the nodes in the network, exhibits a long-tail distribution, implying that while a large number of nodes possess few connections, a small number of nodes contains a large number of connections[37].
+It implies the real neural network have some same statistic properties as social, electrical and transportation networks.
+Also, we plot the correlation curves between degree and cell surface area, volume, and cell length (Fig.1).
+The results indicate that the degree distribution is logarithmically linearly correlated with these characteristics.
+This logarithmic relationship reveals a fundamental principle in the organization of neural networks in the brain.
+To elaborate, as the surface area, volume, or length of a neuron increases, the number of synapses (represented by the degree) it forms grows logarithmically rather than linearly.
+This non-linear scaling is crucial for maintaining efficient information processing in the brain while preventing an explosion in connectivity that would be computationally and energetically costly.
+These characteristics of the brain can explain Weber-Fechner’s law[47]: for any sensory pattern, the perceived intensity is a logarithmic function of the physical intensity.
+Specifically, the size, length, and area changes can only cause a logarithmic change in the number of synapses the neuron has.
+This means that as a stimulus becomes stronger, the increase in neural activity or connectivity is proportionally smaller, mirroring the logarithmic perception of intensity described by Weber-Fechner’s law.
+(a) Visualization of Drosophila Connectome
+(b) Correlation between degree distribution and neuron characteristics
+Figure 1: Visualization and statistic properties of Drosophila connectome data set
+These statistic results are a valuable supplement to previous findings[48] and are crucial for for improved understanding of the Drosophila brain connectcome structure.
+2.3 Activation Patterns under Different Type of Stimuli
+Primarily, to validate the efficacy of our model in signal transmission between input and output neurons, we showcase the activation patterns of intermediate and output neurons throughout the information transmission process.
+This demonstration serves to substantiate that our model accurately reflects real-world scenarios, shaping a quasi-real activation pattern.
+The specific idea for the experiment is following: we stimulate different types of neurons (such as visual input neurons, olfactory neurons, etc.), and record the rate that neurons that are activated in different brain areas.
+If a specific stimulus causes a significantly greater response than the average in the brain area that it is theoretically supposed to activate, then it is considered that the model can generate a quasi-real activation pattern in the brain.
+We give 2 types of input neurons (visual, olfactory) stimulus and record the activation patterns of 6 type of neurons: optic (visual), visual_projection (visual), ALLN (olfactory), ALIN (olfactory), ALPN (olfactory) and ALON (olfactory).
+When we give neuron a stimulus, we set the state of neuron as 1.
+To demonstrate the importance of network structure, we add a comparative model that randomly reconnects networks.
+Specifically, we perform degree-preserving rewiring on a certain proportion of edges in the brain, which means perturbing the network structure while ensuring that the degree of each node remains unchanged.
+In Fig.2 we set reconnect rate $p=10\%$.
+For its sensitivity analysis, please read appendix S4.
+Fig.2 shows the activation patterns of different models.
+From Fig.
+2(a), (e), and (i), it shows visual stimulus cause a strong impact (significantly more than average) on visual neurons (optic, visual projection) while causing a slight impact (significantly less than average) on other type of neurons under all three kinds of models.
+Specifically, under visual stimulus, in response to visual stimuli, the optic (vision related area) exhibit a 36.39%, 4.69% and 9.69% ’s activate rates, respectively.
+At the mean time, no olfactory-related neurons are activated.
+In addition, olfactory stimulus cause a strong impact (significantly bigger than average level) on olfactory neurons (ALLN, ALIN, ALPN, ALON), reaching the activation rate by 94.07%, 78.01%, and 69.75%, respectively.
+And it cause a slight impact (significantly less than average) on other type of neurons.
+For example, it only activates 0.04% of optic and 0.79% of visual_projection in threshold model (Fig.2(c)).
+Although the activation rates of the three models in specific areas differ (due to different activation mechanisms), the significant difference in the activation ratio between correlated and uncorrelated regions indicates the emergence of similar quasi-real activation patterns across all models.
+This suggests that different activation mechanisms are not the fundamental reason for changes in network functionality.
+As for the disturbing models, in Fig.2 (b), (d), (f), (h), (j) and (l), the quasi-real activation patterns totally disappear.
+Across all activation modes and all types of stimuli, no significant differences in activation regions were observed.
+For example, in the threshold model, the maximum difference in the activation ratios of all areas under visual stimuli was only 3.162%.
+Moreover, the range of the regional activation ratios versus the average activation ratio was only 6.32%.
+Additionally, in the sigmoid and LIF models, no significant activation was observed in any area, indicating that stimuli under network disturbance did not lead to any significant response in any area.
+These results indicate that no matter neuron model is, it will lead to similar activation patterns which correspond to reality.
+After disturbing the network structure, the original quasi-real activation patterns in all models completely disappear.
+It means the real network structure matters than activation mechanisms in generating the quasi-real activation patterns.
+Different activation mechanisms are not the primary cause of this realistic activation mechanism; it is the brain’s unique network structure that plays the crucial role.
+This aligns with the studies of complex systems: individual capabilities are not the determining factors in complex systems; what matters is the organization and interaction structure among individuals.
+This perspective highlights the importance of individual relationships and the overall system structure in determining emergent phenomena.
+(a) threhold, visual
+(b) threhold, visual (after disturb)
+(c) threhold, olfactory
+(d) threhold, olfactory (after disturb)
+(e) sigmoid, visual
+(f) sigmoid, visual (after disturb)
+(g) sigmoid, olfactory
+(h) sigmoid, olfactory (after disturb)
+(i) LIF, visual
+(j) LIF, visual (after disturb)
+(k) LIF, olfactory
+(l) LIF, olfactory (after disturb)
+Figure 2: Experiment of intrinsic neurons’ activation. For disturbing model, we conduct 5 repeated experiments and take the average and standard deviation.
+2.4 Distances between Input and Activation Areas
+To further investigate how the network structure, rather than the spatial distribution of neurons, or the parameters of the network model or activation mechanisms, predominantly determines the dynamics of brain activity, we calculated the average network distance and spatial distance between different areas of the brain network.
+The optic and visual_projection area are related to vision and ALLN, ALIN, ALON, ALPN, LHCENT and LHLN are related to olfactory.
+The results are presented in Tab.
+2.
+From Tab.
+2, it is evident that neurons which are closer in spatial distance are not necessarily easier to activate.
+For example, all areas and visual input neurons have similar average spatial distances, but the quasi-real activation patterns can only be observed in specific neural types (optic, visual_projection).
+In contrast, the average network distance between input neurons and neurons should be activated is closer than irrelevant neurons.
+For instance, the distance between visual input neurons and optic, visual_projection are 5.7018 and 5.7432, respectively.
+And the distance between visual input neurons and ALLN, ALIN, ALON, ALPN, LHCENT, LHLN which are irrelevant with vision are 6.6442, 6.2350, 6.1506, 6.5380, 6.3094 and 6.8206, respectively.
+The network distance of neurons related to visual input neurons is generally closer than the distances of unrelated neurons.
+Besides, The average network distance of neurons related to olfactory input neurons is approximately twice as close as the distance of unrelated neurons.
+These results indicates that compared to spatial distance, the average network distance between input neurons and their corresponding activation areas is closer, meaning that compared to spatial structure, network structure is more important, determining these quasi-real activation patterns of the brain network.
+This results reflect the importance of network structure in neural networks, that is, the ability to reproduce certain functions of the brain by considering the network structure only.
+Table 2: Spatial and network distance between input and main areas. The unit of spatial distance is nm and the unit of network distance is the number of edges in the shortest path.
+distance type Neuron type Area Name optic visual_projection ALLN ALIN
+Spatial Distance Visual Input Neurons $365012$ $354725$ $351919$ $348901$
+Olfactory Input Neurons $302690$ $251144$ $110827$ $173345$
+ALON ALPN LHCENT LHLN
+Visual Input Neurons $346925$ $348517$ $362825$ $366013$
+Olfactory Input Neurons $123179$ $115089$ $212515$ $246719$
+optic visual_projection ALLN ALIN
+Network Distance Visual Input Neurons $5.7018$ $5.7432$ $6.6442$ $6.2350$
+Olfactory Input Neurons $5.7967$ $5.1706$ $3.4458$ $3.3845$
+ALON ALPN LHCENT LHLN
+Visual Input Neurons $6.1506$ $6.5380$ $6.3094$ $6.8206$
+Olfactory Input Neurons $3.6312$ $3.6820$ $3.9557$ $3.9758$
+3 Discussion
+This study, through the construction of a large-scale network communication model based on the Drosophila connectome, reveals the critical role of neural network structure in generating brain-like activation patterns.
+Our findings suggest that models based on comprehensive neuron network data can approach the brain’s quasi-real activation patterns even with simple artificial rules.
+This paves the way for understanding biological intelligence, indicating that the structure of biological neural systems may offer valuable reference for designing efficient, fast-adapting intelligent systems.
+Our research emphasizes the importance of network structure, compared to spatial structure or neuronal activation mechanisms, in determining brain activity patterns.
+This finding has significant implications for artificial neural network research, suggesting that future AI designs may need to more closely mimic real brain network structures.
+Furthermore, the large spatial network visualization software we developed provides a new tool for neural network research, facilitating intuitive observation and analysis of complex neural network dynamics.
+Future research directions may include optimizing propagation rules and automatically adjusting link weights to achieve complete simulation of brain behavior.
+These advancements will pave the way for achieving genuine artificial intelligence while deepening our understanding of the nature of biological intelligence.
+4 Acknowledgement
+We thank M.Murthy et al.
+for their invaluable help in sharing the Drosophila connectome data.
+This work was supported by the National Natural Science Foundation of China (72025405, 72088101), the National Social Science Foundation of China (22ZDA102), the Hunan Science and Technology Plan Project (2020TP1013, 2020JJ4673, 2023JJ40685), the Innovation Team Project of Colleges in Guangdong Province (2020KCXTD040), and the Postgraduate Scientific Research Innovation Project of Hunan Province (XJCX2023095).
+The authors declare that they have no conflict of interest.
+APPENDIX
+Appendix A Related Works
 Various models have been developed to mimic or understand aspects of brain function, ranging from the cellular level to the whole brain.
-In this section we review the models that aim to simulate the brain’s behavior.
+In this section, we review the models that aim to simulate the brain’s behavior.
 The review includes neural dynamic model that describes how action potentials in neurons are initiated and propagated.
@@
 At last, we introduce the network communication models which refers to a computational or theoretical framework designed to understand and simulate how different parts of the brain communicate with each other.
-2.1 Neuron Dynamic Model
+A.1 Neuron Dynamic Model
+Neuronal dynamic model is a mathematical framework designed to understand and simulate the electrical activity and dynamic behavior of neurons.
+These models aim to capture the fundamental mechanisms by which neurons process and transmit electrical signals through their membranes, often by using differential equations and computational techniques.
 In past years, neuroscientists have aimed to describing the behavior of neurons by dynamic model.
@@
 The currents are given by:
-$I_{Na}$ $=\bar{g}{Na}m^{3}h(V-E{Na})$ (1)
-$I_{K}$ $=\bar{g}_{K}n^{4}(V-E_{K})$ (2)
-$I_{L}$ $=\bar{g}_{L}(V-E_{L})$ (3)
+$I_{Na}$ $=\bar{g}_{Na}m^{3}h(V-E_{Na}),$ (1)
+$I_{K}$ $=\bar{g}_{K}n^{4}(V-E_{K}),$ (2)
+$I_{L}$ $=\bar{g}_{L}(V-E_{L}),$ (3)
 where $\bar{g}_{Na}$, $\bar{g}_{K}$ and $\bar{g}_{L}$ are the maximum conductances for sodium, potassium, and leak channels.
@@
 The gating variables obey first-order kinetics and are described by the following differential equations:
-$\frac{dm}{dt}$ $=\alpha_{m}(1-m)-\beta_{m}m$ (4)
-$\frac{dn}{dt}$ $=\alpha_{n}(1-n)-\beta_{n}n$ (5)
-$\frac{dh}{dt}$ $=\alpha_{h}(1-h)-\beta_{h}h$ (6)
+$\frac{dm}{dt}$ $=\alpha_{m}(1-m)-\beta_{m}m,$ (4)
+$\frac{dn}{dt}$ $=\alpha_{n}(1-n)-\beta_{n}n,$ (5)
+$\frac{dh}{dt}$ $=\alpha_{h}(1-h)-\beta_{h}h,$ (6)
 where $m$, $n$, and $h$ are the gating variables representing the activation and inactivation states of the ion channels.
@@
 HH model accurately describes the electrophysiological processes of neuronal activity through mathematical equations, allowing scientists to understand the fundamental mechanisms of a group of neurons’ behavior at the molecular and cellular levels.
-Another famous theory is cable theory[16].
+Another classic theory is cable theory[19].
 Cable theory is a mathematical model used to understand how electrical signals propagate along neurons, especially in their dendrites and axons.
@@
 The core equation of cable theory is derived from the cable equation, which is a form of the second-order linear partial differential equation. It can be expressed as:
-$\frac{\partial^{2}V}{\partial x^{2}}=\frac{1}{r_{m}}\left(\frac{\partial V}{\partial t}+\frac{V}{r_{c}}\right)$ (7)
-, where $V$ is the membrane potential, $x$ is the distance along the cable, $t$ is time, $r_{m}$ is the membrane resistance per unit length, $r_{c}$ is the cytoplasmic (axial) resistance per unit length.
-This equation describes how the voltage $V$ changes along the length of the neuron $x$ and over time $t$.
+$\frac{\partial^{2}V}{\partial x^{2}}=\frac{1}{r_{m}}\left(\frac{\partial V}{\partial t}+\frac{V}{r_{c}}\right),$ (7)
+where $V$ is the membrane potential, $x$ is the distance along the cable, $t$ is time, $r_{m}$ is the membrane resistance per unit length, $r_{c}$ is the cytoplasmic (axial) resistance per unit length.
+This equation describes how the voltage $V$ changes along the length of the neuron $x$ and time $t$.
 Cable theory has several important applications in neuroscience, including understanding signal decay, neural modeling and synaptic integration.
 Despite its usefulness, cable theory has limitations.
-First, real neurons have complex geometries, which are simplified in cable theory.
+First, real neurons have complex geometries.
+Cable theory don’t describe well of these complex geometries.
 This can lead to inaccuracies in predicting real neuronal behavior.
-The complex structure of neuron should not be neglected.
 Besides, cable theory primarily addresses passive electrical properties and does not account for active properties like action potentials, which are crucial for neural communication.
@@
 The LIF model captures the essential characteristics of spike generation in neurons.
-The membrane potential $V(t)$ of a neuron at time $t$ is described by the following differential equation.
-$\tau_{m}\frac{dV(t)}{dt}=-[V(t)-V_{rest}]+RI(t)$ (8)
+The membrane potential $V(t)$ of a neuron at time $t$ is described by the following differential equation:
+$\tau_{m}\frac{dV(t)}{dt}=-[V(t)-V_{rest}]+RI(t),$ (8)
 where $\tau_{m}$ is the membrane time constant, which determines how fast the membrane potential decays towards the resting potential $V_{rest}$ in the absence of input.
@@
 Despite their advances, SNNs face significant challenges.
-One primary issue is that the SNN’s structure often draws inspiration from anatomically defined from functional brain networks rather than accurately replicating the intricate neural connections observed at the microscale.
+One primary issue is that the SNN’s structure often draws inspiration from functional brain networks rather than accurately replicating the intricate neural connections observed at the microscale.
 In terms of the connection method of artificial neurons, SNNs usually use a structure similar to that of ANNs.
@@
 These limitations underscore the difficulties in considering SNNs as the optimal approach for brain simulation in their current state.
-As for the large scale brain simulation model aims to simulate the whole brain’s behavior, Eugene et al.[25] present the first large-scale brain simulation model based on experimental measures in several mammalian species.
+As for the large scale brain simulation model aims to simulate the whole brain’s behavior, Eugene et al.[28] presents the first large scale brain simulation model based on experimental measures in several mammalian species.
 The model is founded upon the global thalamocortical anatomy derived from diffusion tensor imaging of the human brain.
-It incorporates multiple thalamic nuclei and a six-layered cortical microcircuitry, which is grounded in in vitro labeling and three-dimensional reconstruction of single neurons from the cat visual cortex.
-Additionally, the model simulates over 1,000,000 SNN neurons, calibrated to replicate known in vitro response patterns recorded in rats.
+It incorporates multiple thalamic nuclei and a six-layered cortical microcircuitry, which is grounded in three-dimensional reconstruction of single neurons from the cat visual cortex.
 It boasts nearly half a billion synapses, incorporating accurate receptor kinetics, short-term plasticity, and dendritic spike-timing-dependent synaptic plasticity (dendritic STDP).
@@
 This model is a comprehensive attempt to simulate the whole neural system, which is a critical component of the brain involved in sensory processing and consciousness.
-Eliasmith et al.[26] introduce a more comprehensive brain simulating model, which consists of 2,500,000 neurons and serves to close this divide by demonstrating a variety of behaviors.
+Eliasmith et al.[29] introduce a more comprehensive brain simulating model, which consists of 2,500,000 SNN neurons.
 This model receives only sequences of visual images and produces all its responses through a simulated arm.
@@
 As for the standardized platform, Zeng et al.
-[27] present BrainCog, a cognitive intelligence engine inspired by the brain.
+[30] presents BrainCog, a cognitive intelligence engine inspired by the brain.
 This platform, based on SNN, offers critical infrastructure for the creation of brain-inspired artificial intelligence and simulation of the brain.
 BrainCog encompasses a variety of biological neurons, encoding methods, learning principles, brain regions, and a combination of hardware and software as its core elements.
-However, it is essential to note that each neuron’s dynamic model involves multiple differential equations, posing a significant challenge in simulating the complete behavior of the intact brain given the constraints of available computing power.
-Therefore, there is a compelling need for a cost-effective simulation approach to address this challenge.
+However, large scale brain simulations require the simulation of millions or even billions of neurons, posing extremely high demands on computational resources.
+Even the most powerful supercomputers might need weeks or even months to complete a single simulation.
+Additionally, as the number of neuron connections and synaptic weights increases, data storage and processing requirements also significantly rise.
+Besides, although large scale brain simulations aim to simulate neurons and their connections, they often fail to capture all the biological details present in the brain.
+Factors like neuron morphology, molecular-level dynamics, and changes in gene expression are complex biological phenomena that are difficult to fully integrate into the models.
+Thus, the simulation results might not fully reflect the real behavior of the brain.
+At last, building a large-scale brain simulation model requires a vast amount of precise biological data, such as the connectivity between neurons, synaptic strengths, types, and concentrations of neurotransmitters.
+However, acquiring these data is often very challenging, involving complex and expensive experimental techniques, and it’s difficult to measure comprehensively across the entire brain.
 As for the collective action of neuron populations, despite the success of neural system mean field models[31] in areas such as modeling epileptic seizures[32], there is still a lack of models that precisely describe the large-scale activity of whole brain.
-2.2 Brain Network Communication Model
-The network communication model is a conceptual framework that describes how data is transmitted from one node to another over a network.
-It breaks down the complex process of communication into simpler, more manageable layers or components, each with a specific role.
+A.2 Brain Network Communication Model
+The network communication model is a conceptual framework which describes how data is transmitted from one node to another over a network.
+It breaks down the complex process of communication into simpler, more manageable layers or components.
 The most widely recognized network communication models are the Open Systems Interconnection (OSI) model and the Internet Protocol Suite, commonly known as the TCP/IP model.
-Nowadays, scientists apply network communication models to neuron research, which are mostly used to understand how brain networks operate[30][31][32].
+Nowadays, scientists apply network communication models to neuron research, which are mostly used to understand how brain networks operate[33, 34, 35].
 Early studies have indicated that, similar to social and electrical networks[36], brain networks also exhibit properties such as power-law distributions[37] and the rich-club phenomenon[38].
 These results indicate that brain networks contain a wealth of network statistical characteristics.
-To reveal the mystery of the brain, researchers must take network characteristics into consideration.
 Researchers designed many brain network models to describe the processes of message propagation in brain.
@@
 The most famous model of this category is shortest path model[39], which assumes that signals between a pair of nodes travel along the shortest path between them.
-Let $d(v)$ represent the shortest path length from the starting node to node $v$, for each node $v$ in the network, initialize $d(v)=\infty$, then for the starting node $s$, set $d(s)=0$.
-Next, process each node in increasing order of $d(v)$, for each node, update the $d(u)$ value of its neighbors $u$.
+In the shortest path model, the shortest path length from the starting node $s$ to each node $v$, is initialized as $d(s)=0$ and $d(v)=\infty$; then, process each node in increasing order of $d(v)$, for each node, update the $d(u)$ value of its neighbors $u$.
 The update way can be represented as follows:
-$d(u)=\min(d(u),d(v)+w(v,u))$ (9)
+$d(u)=\min(d(u),d(v)+w(v,u)),$ (9)
 where $w(v,u)$ is the weight of edge $e_{uv}$.
+Information is propagated entirely through the shortest paths in these models, which means in terms of propagation cost, signal transmission is optimal.
 However, the key disadvantage of the model is these methods require each neuron must have knowledge about the entire network structure[40], which doesn’t correspond to the reality: every neuron only has the message of its neighbor neurons.
-[…]
+Another well-known model is navigation model.
+It suggests that neurons always send information to the neuron that can be reached with the least communication cost[41].
+These models achieve the lowest communication cost.
+A significant achievement is that some algorithms can identify paths with 70%-100% of optimal signaling efficiency in the functional connectomes of different organisms[42].
 However, the main disadvantage of navigation model is it supposes that neurons know the distance between their neighbors and a desired target region.
@@
 For unbiased random walk, the function of random walk model is as follows.
-$P_{ij}=\frac{1}{k_{i}}$ (10)
+$P_{ij}=\frac{1}{k_{i}},$ (10)
 where $P_{ij}$ represent the probability of node $i$ choosing node $j$.$k_{i}$is the degree of node $i$.
 As for biased random walk, for example, if the transition probability depends on the edge weights or certain properties of the nodes, then the transition probability can be adjusted to
-$P_{ij}=\frac{w_{ij}}{\sum_{l\in N(i)}w_{il}}$ (11)
+$P_{ij}=\frac{w_{ij}}{\sum_{l\in N(i)}w_{il}},$ (11)
 where $w_{ij}$ is the weight of node $i$ to node $j$,$N(i)$is the set of node $i$.
-However, studies have shown that the efficiency of signal transmission through random walks is relatively low[41]. So it can’t describe the rapid communication process in the brain.
+The diffusion process does not require individual neurons to have knowledge about the connectivity of connectome beyond their immediate neighbors.
+However, studies have shown that the efficiency of signal transmission through random walks is relatively low[44].
+So it can’t describe the rapid communication process in the brain.
 An effective category is threshold model[45].
@@
 The condition for node $i$ to adopt the behavior at time $t$+1 can be formulated as follows.
-$a_{i}(t+1)=1\quad\text{if}\quad\frac{1}{k_{i}}\sum_{j\in N_{i}}a_{j}(t)\geq\theta_{i}$ (12)
+$a_{i}(t+1)=1\quad\text{if}\quad\frac{1}{k_{i}}\sum_{j\in N_{i}}a_{j}(t)\geq\theta_{i},$ (12)
 where $N_{i}$ is the set of neighbors of node $i$, $a_{i}(t)$ is the adoption state of node $i$ at time $t$.
@@
 $k_{i}$ is the degree of node $i$.
-The threshold model has been extensively applied in artificial neural networks.
+The threshold model has stronger biological interpretability because it approximates the mechanism of neuronal firing.
+Consequently, the threshold model has been extensively applied in artificial neural networks.
 However, how to set the propagation thresholds for different neurons, and how to design the dynamics model between neurons remains a challenge.
-3 Data and Methods
-3.1 Drosophila Connectome Data Set
+Appendix B Data and Methods
+B.1 Drosophila Connectome Data Set
 We employ the recently released largest Drosophila brain connectome[17], offering a thorough mapping of the entire neural network within a Drosophila female brain.
@@
 For a more in-depth understanding, please refer to [17][46] for detailed information.
-Table 1 shows the basic properties of the data set.
+In this section, we use nodes to represent neurons and connections to represent the synapses between neurons.
+Table 3 presents the basic properties of the data set.
 Note that we combine multiple edges between two nodes to calculate network statistic properties.
-The average degree before combining multiple edges is 501.6 while average degree is 37.01 after combination.
-It means most of the pairs of neuron has many links.
+The average degree before and after combining multiple edges is 501.6 and 37.01, respectively, indicating it is prevalent that the pairs of neurons are connected by multiple synapses.
 Table 3: Statistical properties of Drosophila Brain Connectcome Graph
@@
 $501.6$ $0.1527$ $0.0008817$
-Apart from that, we draw the curves of degree distribution.
-From Fig.
-1 We found that the distribution of log(degree) initially appears to be approximately linear, but at the end it exhibits a long-tail distribution, which demonstrate that the original distribution of degree follows a power-law distribution.
-This result corresponds to the earlier research that neural network also has the power-law property[34].
-Figure 1: Degree Distribution (Log-Log Scale)
-Also, we plot the correlation curves between degree and cell surface area, volume, and cell length (See in Fig.
-2).
-The results indicate the degree distribution is logarithmically linearly correlated with these data.
-These characteristics of the brain can explain Weber-Fechner’s law[44]: for any sensory pattern, the perceived intensity is a logarithmic function of the physical intensity.
-The structural changes can only cause a logarithmic change in the average connection strength.
-Figure 2: Correlation between degree distribution and neuron characteristics
-These statistic results are a valuable supplement to previous findings[45] and are crucial for a deeper understanding of this data set.
-3.2 Model
+Network Diameter Assortativity Coefficient
+$11$ $0.0008817$
+B.2 Model
 In this section, we introduce the large scale network communication model which consumes less computational resource and can efficiently utilize the network structure.
-First we introduce the symbol systems in this article.
+To solve the problem that the link weights between neurons can’t be measured, we propose two hypotheses to estimate the link weights.
+In addition, we replace the node model in the network communication model with an improved LIF neuron model for comparison.
+B.2.1 Symbol Systems
 The original network presentation of neural systems often uses nodes to represent the neurons and using edges to represent the links (chemical synapses, electrical synapses, etc) between neurons[4].
-However, this approach has a number of limitations: Firstly, there are more than one link (synapse) between two neurons[46] and the links are spatially and temporally asymmetric.
-To illustrate, two neurons shown in Fig.
-3 have multiple synapses between them, and the dynamical properties of these different synapses may vary.
-Additionally, due to their spatial locations, the signal transmission delay and the position of action on the downstream neurons are also different, making it difficult to simply merge them.
-Figure 3: An example of Complex Connection
-For this reason, similar to previous research[47], we use the dual method on Drosophila connectome data set. The method is introduced below.
-A dual graph is a concept in graph theory which describes a special relationship between graphs.
+However, this approach has a number of limitations: first, there are usually more than one link (synapse) between two neurons[49] and the links are spatially and temporally asymmetric.
+To illustrate, two neurons shown in Fig.S1 have multiple synapses between them, and the dynamical properties of these different synapses may vary.
+Additionally, due to their spatial locations, the signal transmission delay and the position of action on the downstream neurons are also different, making it an oversimplification if the synapses are merged into one edge.
+Figure S1: An example of complex connection
+For this reason, we use the dual method[50] on Drosophila connectome data set to model the multi-link between neurons.
 In a dual graph, the nodes and edges of the original graph are interchanged, resulting in a new graph.
@@
 Fig.
-4 is an example of an original graph and a dual graph.
-After being processed, the new graph can be considered to be a hyper graph, whose nodes represent the synapses and links represent the cell body.
-Then we can use the hyper communication model to describe it.
+S2 is an example of an original graph and a dual graph.
+After being processed, the nodes in the dualed graph represent the synapses and links represent the cell body.
+Then we can use the communication model to describe the message propagation process.
 (a) original graph
 (b) dual graph
-Figure 4: An example of original graph and dual graph
-In order to standardize the symbol system, we use a set to represent the neuronal cell body and use an element to represent a physical link between neurons.
-We regard the process of messages interactions as the elements’ interactions on sets.
-Let $N=\{n_{1},...,n_{i},...\}$ represent the set of all elements and $E=\{e_{1},...,e_{j}...\}$ represent the probable sets of elements.
-Note that the intersection of sets include some elements, which is formalized as the following manner.
-$e_{k}\cap e_{q}=\{n_{i}|(n_{i}\in e_{k})\wedge(n_{i}\in e_{q})\}$ (13)
-Note that union of every element in $N$ and $E$ is the whole graph. Let $G$ represent the graph. The graph’s definition is given as follows.
+Figure S2: An example of original graph and dual graph
+In order to standardize the symbol system, we use a edge (multi-edge, could include more than 2 nodes) to represent the neuronal cell body and use an node to represent a physical link between neurons (synapse).
+We regard the process of messages interactions as the nodes’ interactions via edges.
+Let $N=\{n_{1},...,n_{i},...\}$ represents the set of all nodes and $E=\{e_{1},...,e_{j}...\}$ represents the sets of all edges. An obvious relationship is formalized as the following manner.
+$e_{k}\cap e_{q}=\{n_{i}|(n_{i}\in e_{k})\wedge(n_{i}\in e_{q})\}.$ (13)
+Note that the union of each element in $N$ and $E$ is the whole graph. The graph $G$’s definition is given as follows.
 $G=\bigcup_{j}e_{j}=\bigcup_{i}\{n_{i}\}.$ (14)
-Because neural link is directed, let ${e_{j}}^{+}$ ( ${e_{j}}^{-}$) represent the set of elements which output (input) the signals, note that ${e_{j}}^{+}$ = ${e_{j}}^{+}\cup{e_{j}}^{-}$.
-Every neuron and synapse has a activate state (state = 1) or resting state (state = 0).
+Let the label of any neuron be $j$.
+Because neural link is directed, let ${e_{j}}^{+}$ ( ${e_{j}}^{-}$) represent the set of elements which output (input) the signals, note that ${e_{j}}$ = ${e_{j}}^{+}\cup{e_{j}}^{-}$.
+Each neuron and synapse has an activate state (state = 1) or resting state (state = 0).
+When a neuron is at the activate state, it fires to other neurons, otherwise it keeps rest.
 So set the statement of the element $i$ as $X_{i}$.
@@
 $Y_{j}=1$ if the set $e_{j}$ is active, otherwise $Y_{j}=0$.
-We use a linear threshold activation function. Then, general transmission function is defined in the following manner,
-$Y_{j}=\sum_{\{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset\}}\sum_{n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\sigma(\omega(n_{k},e_{i},e_{j})X_{k})$ (15)
-where $\omega(n_{k},e_{i},e_{j})$ represents the link weight between set $i$ and set $j$ (specifically, two neurons).
+B.2.2 Large Scale Network Communication Model
+By modeling the Drosophila connectome as graph model, we introduce the large scale network communication model on this graph.
+We use a linear threshold activation function on each neuron.
+Thus, general transmission function is defined in the following manner,
+$Y_{j}=\sigma(\frac{1}{\beta_{j}}\sum_{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset}\sum_{k|n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\omega(n_{k},e_{i},e_{j})X_{k}),$ (15)
+where $\omega(n_{k},e_{i},e_{j})$ represents the link weight between set $i$ and set $j$ (specifically, two neurons) through element $n_{k}$.
 Here we consider that the link weight between two neurons is not only related to the neurons themselves, but also to the synapses between them.
 $\sigma(.)$ is the activate function.
+$\beta_{j}$ is the normalization factor.
+In this paper we use threshold activation function and sigmoid function, which are two .
+The threshold activation function is a simple, binary decision function used in artificial neurons.
+It outputs a value of 1 if the input exceeds a certain threshold, and 0 otherwise.
+This function is also known as the Heaviside step function.
+$\sigma(x)=\left\{\begin{aligned} 1\quad x\geq threshold\\ 0\quad x<threshold\\ \end{aligned}\right.$ (16)
+The sigmoid activation function is a smooth, S-shaped function that maps input values to an output range between 0 and 1.
+It is commonly used in neural networks to introduce non-linearity and to model probability-like outputs in binary classification problems.
+The sigmoid function is defined as the logistic function.
+$\sigma(x)=\frac{1}{1+e^{-x}}$ (17)
 The statement of an neuron will influence the statement of all output synapses. The equation between neuron $j$ and output synapse $i$ is as the following manner.
-$X_{i}=Y_{j}$ (16)
+$X_{i}=Y_{j}\ \ \ \ if\ n_{i}\in e^{+}_{j}.$ (18)
 The link weights between neurons determine the extent of influence they have on each other.
@@
 However, existing experimental methods are not well-suited for directly measuring the weights between neurons in a whole brain, so adopting some alternative approach is necessary.
-Below, we introduce two hypotheses that use different methods to estimate the connection weights between neurons.
-3.2.1 Average Hypothesis
+Below, we introduce a hypothesis that use different methods to estimate the connection weights between neurons.
 As previously mentioned, there exists a large number of synapses between two neurons.
@@
 That is to say,
-$\omega(n_{k},e_{i},e_{j})=1$ (17)
+$\omega(n_{k},e_{i},e_{j})=1.$ (19)
+Additionally, We denote the sum of all link weights between neurons $i$ and $j$ as $W_{ij}$:
+$W_{ij}=\sum_{k|n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\omega(n_{k},e_{i},e_{j}).$ (20)
+And we define $\beta_{j}$ as the maximum weight among all $W_{ij}$ related to neuron $j$. That is to say,
+$\beta_{j}={{Max}(\{W_{ij}|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset}\}).$ (21)
+Where $Max(S)$ represents the Maximal element in set $S$.
 Then the normalized equation is as follows.
-$Y_{j}=\sigma(\frac{\sum_{\{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset\}}\sum_{n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\sigma(\omega(n_{k},e_{i},e_{j})X_{k})}{{{Max}(\{\omega(n_{k},e_{i},e_{j})|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset,n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\})})$ (18)
-3.2.2 Distance Hypothesis
-The continuous evolution of organisms has led to the brain’s spatial structure being highly complex, with a small area possibly containing a large number of neurons.
-Besides, a fact is messages are transmitted through bioelectric currents.
-Assuming that the distance is longer, the resistance will be greater, resulting in a smaller communication effect.
-Based on this, there emerges a hypothesis: Is the strength of interactions between neurons inversely proportional to the spatial distance between them?
-So in the distance hypothesis we assume that the impact of input synapses on neurons is inversely proportional to the distance between them.
-Then the norm can be formulated as follows.
-$\omega(n_{k},e_{i},e_{j})=(\parallel\vec{r_{k}}\parallel^{2}-\parallel\vec{r_{j}}\parallel^{2})^{\frac{1}{2}}$ (19)
-Where ${r_{k}}$, ${r_{j}}$ are the center position vector of synapse and cell body, respectively.
-Then the transmission function is defined as the following manner.
-$Y_{j}=\sigma(\frac{\sum_{\{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset\}}\sum_{n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\sigma(\omega(n_{k},e_{i},e_{j})X_{k})}{{{Max}(\{\omega(n_{k},e_{i},e_{j})|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset,n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\})})$ (20)
-3.3 Evaluation Criteria
+$Y_{j}=\sigma(\frac{1}{\beta_{j}}\sum_{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset}\sum_{k|n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\omega(n_{k},e_{i},e_{j})X_{k}).$ (22)
+B.2.3 Neuron Dynamic Model
+To better reflect real-world conditions and demonstrate that the superiority of this network architecture does not solely rely on the design of the activation function, we also employed a variant of the LIF model[23] as the basic neuron model.
+The membrane potential change of a neuron’s voltage can be described by the following differential equation.
+$V_{j}(t+dt)=V_{j}(t)+\left(\frac{dt}{\tau}\right)\left(R\cdot I^{syn}_{j}-(V_{j}(t)-V_{\text{reset}})\right)$ (23)
+Where $V_{j}$ represents the membrane potential of $j^{th}$ neuron, $R$ represents the resistance constant and $I^{syn}_{j}$ represents the total input current of neuron $j$.
+$\tau$ represents the time constant.
+The input current $I^{\text{syn}}_{j}$ can’t be calculated directly. To estimate the value of $I^{\text{syn}}$, we make the following assumptions based on the following supposes:
+1. The magnitude of the current on the synapse should be proportional to the voltage difference between the connected neurons.
+2. The current on each synapse should be proportional to the connection weight between the neurons.
+3. A current can only be emitted if the output neuron is in an active state.
+Considering the supposes above, the equation of input current $I^{\text{syn}}_{j}$ is designed as follows:
+$I^{\text{syn}}_{j}=\frac{1}{\alpha\beta_{j}}\sum_{{}_{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset}}\sum_{k|n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\left(\omega(n_{k},e_{i},e_{j})\times(V_{i}-V_{j})\times Y_{i}\right),$ (24)
+where $V_{i}-V_{j}$ represents the voltage difference between neuron $i$ and $j$.
+$Y_{i}$ represents the state of output neuron $i$.
+$\beta_{j}$ is defined according to equation (21).
+$\alpha$ is the proportion const.
+According to the LIF model, when a neuron’s voltage reaches the threshold value, the neuron will fire and then quickly resets its voltage.
+Based on this, the mechanism for updating the neuron voltage is designed as follows.
+$\text{if}\ V_{j}\geq V_{th}&:\ Y_{j}=1\ \text{and}\ V_{j}=V_{\text{rest}}\\ \text{else}&:\ Y_{j}=0\end{aligned},$ (25)
+where $V_{th}$ represents the fire voltage and $V_{\text{rest}}$ represents the rest voltage, respectively.
+Other propagation mechanism is consistent with network model introduced in section B.2.2.
+B.3 Evaluation Criteria
 In previous study, researchers used classic indicators such as signaling cost and activation time to evaluate the quality of the model[16][51][52].
-However, these indicators can’t prove the model accords with the true communication process even if the activation time is short and the signaling cost is minimum, etc.
-To better assess whether a model can reflect the real situation of brain communication, we use the percentage of activated neurons in regions that should work under a certain stimulus.
-To illustrate, if the brain gets a visual stimulus, the region related to vision should work while the region not related to vision should keep in a resting state.
-Then we record the percentage of neurons activated and if regions related to the certain function have a significant higher activation percentage than the average activation percentage of the whole brain, we call it a quasi-real activation pattern.
-We define the consistency indicator $C$. The consistency indicator $C$ in region $r$ which has a certain function (visual, olfactory, etc) is defined in the following manner.
-$C_{r}=N_{r}^{act}/N_{r}$ (21)
+However, these indicators can’t describe the model accords with the true communication process or not, even if the activation time is short and the signaling cost is minimum, etc.
+To better assess whether a model can reflect the real situation of brain communication, we use the percentage of activated neurons in regions that related to (should work under) a certain stimulus.
+To illustrate, if the brain gets a visual stimulus, the region related to vision should work while the region not related to vision should remain a resting state.
+Then we record the percentage of neurons activated.
+If regions related to the certain function have a significant higher activation percentage than the average activation percentage of the whole brain, we call it a quasi-real activation pattern.
+We define the consistency indicator $C_{r}$ in region $r$ which has a certain function (visual, olfactory, etc) is defined in the following manner.
+$C_{r}=N_{r}^{act}/N_{r},$ (26)
 where $N_{r}^{act}$ represents the number of neurons activated in area $r$ at the final state, $N_{r}$ represent the number of neurons in area $r$ and $N$ represent the total number of neurons.
-3.4 Large Spatial Network Visualization
-The visualization software of experiment observation on large spatial networks is important for researchers[50][51].
-However, there is still a lack of visualization system that can observe information propagation on large spatial networks in real-time, which poses difficulties in experiment.
-To observe directly the communication progress on this super large Drosophila connectome, we develop a large spatial network visualization software.
-The drawing part mainly employs the Canvas provided by HTML5 and we use the Three.js based on WebGL technology to achieve spatial network visualization effects, while the user interface display is constructed by Vue.js, providing users with an intuitive and interactive interface to display and analyze data.
-Fig.
-5 is an example of the 3D visualization system.
-This software will be made available to other researchers with the author’s permission.
-(a) Side view 1
-(b) Side view 2
-(c) Side view 3
-(d) Top view 1
-(e) Top view 2
-(f) Top view 3
-(g) Front view 1
-(h) Front view 2
-(i) Front view 3
-Figure 5: Visualization of Spatial Drosophila Brain Network
-3.4.1 Front End Technical Details
-We use the Vue.js framework to build the user interface.
-Vue.js is a popular JavaScript framework that simplifies the development of web applications through componentization, enabling the creation of front-end interfaces with rich interactivity and good user experience.
-The rendering of three-dimensional complex networks relies heavily on the Three.js library.
-Three.js is a powerful JavaScript library for creating and displaying 3D graphics.
-It offers a rich set of 3D objects, materials, lights, and animations, allowing for the easy creation of high-quality 3D scenes and objects.
-To improve rendering performance, GPU acceleration technology is utilized.
-GPU acceleration takes advantage of the computational power of the Graphics Processing, significantly enhancing rendering speed and efficiency through parallel processing and stream processing modes.
-The GPU acceleration capabilities of Three.js, especially through the creation of point objects using Float32BufferAttribute, which contains their position information in three-dimensional space, greatly enhance the rendering speed and smoothness.
-3.4.2 Back End Technical Details
-We use the Spring Boot framework, an open-source framework based on Java for quickly building independent, production-grade Spring applications.
-With Spring Boot, backend services can be rapidly set up to provide RESTful API interfaces for frontend consumption.
-The backend also utilizes the PostgreSQL database to store network data.
-Besides, We use PostgreSQL, which is a powerful open-source relational database management system that offers a rich set of data types and robust querying capabilities.
-To store 3D network data, we leverage PostgreSQL’s JSONB type, which allows storing JSON-formatted data within PostgreSQL and provides a series of functions and operators for querying and manipulating this data.
-With the JSONB type, information about nodes and edges in the network can be stored, enabling fast querying and updating operations.
-4 Experimental Results and Analysis
-In this part we introduce the experimental results and analysis.
-Above all, we introduce the activation pattern experiment.
-We give different stimuli and record the activation pattern of different neurons.
-In addition, we give a unilateral stimulus (stimulus one hemisphere) and record the bilateral response of neurons.
-Finally, to prove the network structure matters in brain network communication, we calculate the average physical and network distance between different areas.
-4.1 Experimental Settings
-The basic methodology of the experiment is described as follows.
-In this study, a continuous stimulus signal (binary 0-1 signal) is applied to a subset of input neurons (such as visual or olfactory neurons) in the brain.
-Subsequently, we compute the signal’s propagation process within the brain and record the responses of both intermediate and output neurons.
-The evaluation criterion is determined by calculating the ratio of neurons that ought to be activated to those that are actually activated in a specific area.
-Given that the brain functions as a type of shallow neural network[52], we document the outcomes at an iteration step size of 5, by which point the majority of neurons have achieved a state of stability.
-Moreover, as the input remains constant and no randomization parameters have been introduced, the outcomes of each experiment are consistent, eliminating the need for repetition.
-Next, we will introduce the background knowledge of the neurons involved in this paper. Here are introductions of these neurons.
+Appendix C Introduction of Related Neurons
 Optic neurons, also known as retinal ganglion cells, are the first neurons in the visual pathway.
@@
 For further information, please read [46].
-4.2 Activation Patterns on Different Type of Neurons
-Primarily, to validate the efficacy of our model in modifying signal transmission between input and output neurons, we showcase the activation patterns of intermediate and output neurons throughout the information transmission process.
-This demonstration serves to substantiate that our model accurately reflects real-world scenarios, shaping a quasi-real activation pattern.
-The specific idea for the experiment is following : we stimulate different types of neurons (such as visual input neurons, olfactory neurons, etc.), and record the rate that neurons that are activated in different brain areas.
-If a specific stimulus causes a significantly greater response than the average in the brain area that it is theoretically supposed to activate, then it is considered that the model can generate a quasi-real activation pattern in the brain.
-We give 2 types of input neurons (visual,olfactory) stimulus display 6 type of neurons: optic (visual), visual_projection (visual), ALLN (olfactory), ALIN (olfactory), ALPN (olfactory) and ALON (olfactory).
-To demonstrate the importance of network structure, we add a comparative model that randomly reconnects networks.
-Specifically, we calculate the average network distance from input neurons (visual/olfactory) to corresponding neurons (optic, visual_projection/ALLN, ALPN, ALON, ALLN), round it to the nearest integer to determine the average layer of that neuron, and then randomly swap the connections between nodes within each layer.
-Fig.
-6 shows the activation patterns of different models.
-The different neuron types are on the horizontal ordinate and the activation percentage is on the vertical ordinate.
-The horizontal line represents the percentage of activated neurons in the whole brain.
-From Fig.
-6 we can find that under the average hypothesis, visual stimulus cause a strong impact (significantly more than average) on visual neurons (optic, visual projection) while causing a slight impact (significantly less than average) on other type of neurons.
-Specifically, under visual stimulus, in response to visual stimuli, the optic (vision related area) exhibit a 10.26% higher activation rate compared to the average level, while no olfactory-related neurons are activated.
-In addition, olfactory stimulus cause a strong impact (significantly bigger than average level) on olfactory neurons (ALLN, ALIN, ALPN, ALON), exceeding the average level by 81.49%.
-And it cause a slight impact (significantly less than average) on other type of neurons.
-For example, it only activates 0.04% of optic and 0.79% of visual_projection.
-Furthermore, the results are not as significant on the distance model.
-Some incorrect areas were activated: The activate rate of ALLN is only 7.08%, with a 5.16% difference from the average level (Fig.
-6(c)).
-Besides, ALIN, ALPN and ALLN, which should not be activated, have activate rates of 12.50%, 0.15%, and 1.89%, respectively (Fig.
-6(c)).
-These results indicate that whether distance hypothesis can describe the actual situation remains to be discussed, which is a refutation for previous research[53].
-As for the shuffle model, the ALIN, ALPN, ALON and ALLN which should not be activated have activate rates of 70.00%, 54.13%, 61.25% and 62.31%, respectively (Fig.
-6(e)).
-Although these activate rates do not meet the average level, they are significant activation rates that means many neruons are activated incorrectly.
-Moreover, optic and visual_projection are activated incorrectly, with activate rates of 55.11% and 42.52%, respectively (Fig.
-6(f)).
-Besides, ALLN, ALPN, ALON, and ALLN which are supposed to exceed the average level are actually lower than the average level by 0.01%, 17.74%, 21.25% and 9.34%, respectively (Fig.
-6(f)).
-These results indicate that after shuffling the network structure, the original quasi-real activation patterns completely disappear.
-It means the real network structure matters in generating the quasi-real activation patterns.
-In addition, even just designing simple models, it is feasible to simulate the propagation of information in the brain at the level of neurons and generate a quasi-real activation pattern.
-(a) visual stimulus, average model
-(b) olfactory stimulus, average model
-(c) visual stimulus, distance model
-(d) olfactory stimulus, distance model
-(e) visual stimulus, shuffle model
-(f) olfactory stimulus, shuffle model
-Figure 6: Experiment of intrinsic neurons’ activation. $\sigma_{0}=0.8$. For shuffle model, we conduct 5 repeated experiments and take the average and standard deviation.
-4.3 Distances between Input and Activation Areas
-In the previous section, the quasi-real activation patterns on different type of neurons of the brain was presented.
-Through observation, this relationship was identified, and a quasi-real activation pattern was observed in the drosophila brain.
-There is an issue here: Is the quasi-real activation pattern determined by the brain network’s structure, the spatial structure of the brain, or the quality of the network communication model?
-Based on the puzzle, we calculated the average network distance and spatial distance between different areas of the brain network.
-The optic and visual_projection area are related to vision and ALLN, ALIN, ALON, ALPN, LHCENT and LHLN are related to olfactory.
-The results are presented in Tab.
-2.
-From Tab.
-2, it is evident that neurons which are closer in spatial distance are not necessarily easier to activate.
-For example, all areas and visual input neurons have similar average spatial distances, but the quasi-real activation patterns can only be observed in specific neural types (optic, visual_projection).
-In contrast, the average network distance between input neurons and neurons should be activated is closer than irrelevant neurons.
-For instance, the distance between visual input neurons and optic, visual_projection are 5.7018 and 5.7432, respectively.
-And the distance between visual input neurons and ALLN, ALIN, ALON, ALPN, LHCENT, LHLN which are irrelevant with vision are 6.6442, 6.2350, 6.1506, 6.5380, 6.3094 and 6.8206, respectively.
-The network distance of neurons related to visual input neurons is generally closer than the distances of unrelated neurons.
-Besides, The average network distance of neurons related to olfactory input neurons is approximately twice as close as the distance of unrelated neurons.
-These results indicates that compared to spatial distance, the average network distance between input neurons and their corresponding activation areas is closer, meaning that compared to spatial structure, network structure is more important, determining these quasi-real activation patterns of the brain network.
-This results reflect the importance of network structure in neural networks, that is, the ability to reproduce certain functions of the brain by considering only the network structure.
-Table 2: Spatial and network distance between input and main areas. The unit of spatial distance is nm and the unit of network distance is the number of edges in the shortest path.
-distance type Neuron type Area Name optic visual_projection ALLN ALIN
-Spatial Distance Visual Input Neurons $365012$ $354725$ $351919$ $348901$
-Olfactory Input Neurons $302690$ $251144$ $110827$ $173345$
-ALON ALPN LHCENT LHLN
-Visual Input Neurons $346925$ $348517$ $362825$ $366013$
-Olfactory Input Neurons $123179$ $115089$ $212515$ $246719$
-optic visual_projection ALLN ALIN
-Network Distance Visual Input Neurons $5.7018$ $5.7432$ $6.6442$ $6.2350$
-Olfactory Input Neurons $5.7967$ $5.1706$ $3.4458$ $3.3845$
-ALON ALPN LHCENT LHLN
-Visual Input Neurons $6.1506$ $6.5380$ $6.3094$ $6.8206$
-Olfactory Input Neurons $3.6312$ $3.6820$ $3.9557$ $3.9758$
-4.4 Bilateral Response under Unilateral Stimulus
+Appendix D Bilateral Response under Unilateral Stimulus
 In real life, the left and right hemispheres cooperate to perform complex tasks.
@@
 More precisely, we independently simulate visual input neurons in the left and right hemispheres and record the rate of activated neurons in each hemisphere over time.
-Figure 7 illustrates the curve depicting how the response ratio evolves with each time step.
-It’s noteworthy that when simulating one hemisphere alone, visual processing neurons in both hemispheres are activated.
+Fig.S3 illustrates the curve depicting how the response ratio evolves with each time step.
+The horizontal line represents the proportion of activated neurons in all hemisphere.
+It is noteworthy that when simulating one hemisphere alone, visual processing neurons in both hemispheres are activated.
 At time step 5, the unstimulated hemisphere begins to have a activate pattern, which corresponds to the hypothesis of shallow connection[53].
@@
 (b) Right stimulus
-Figure 7: Experiment of activation rate under single stimulus. We set threshold $\sigma_{0}=0.8$.
-Visually, the shape of the brain appears to be physically similar, but whether the network structures are similar remains unknown.
-To explore the principle behind similar responses caused by a unilateral stimulus, we calculated the degree, clustering coefficient, and eigenvector centrality distribution for each hemisphere.
-The Pearson correlation coefficient of degree, clustering coefficient, and eigenvector centrality vector is 0.9986, 0.9988, and 0.9991, respectively.
-According to Fig.
-8, these statistical indicators for both hemispheres are extremely similar, indicating that the brain exhibits strong symmetry in network structure.
-(a) Degree
-(b) Clustering
-(c) Eigenvector centrality
-Figure 8: Degree, clustering coefficient, and eigenvector centrality distribution
-5 Conclusion
-Considering the insufficient evidence supporting the vitalness of network structure in brain message transmission and the effectiveness of network models in simulating brain activation patterns, we propose a large scale network communication model based on simple propagation rules.
-We also design an evaluation criteria to compare real and simulated activation patterns.
-Our research utilizes the largest adult Drosophila connectome data set, where we analyze the basic statistical properties and explore the relationship between neuron length, area, size, and connectivity degree.
-Experimental results demonstrate that our model produces an activation pattern across the brain that closely resembles reality when subjected to identical stimulus on input neurons.
-This result strongly suggests that even with simple propagation rules, network models can approximate the brain’s activation patterns.
-Besides, when changing the network structure, there is an incorrect activation pattern on the brain network.
-It is an evidence that network structure matters in generation real brain activation patterns.
-Furthermore, to prove network structure matters rather than propagation rules or spatial structure, we assess the network and spatial distances across different areas, discovering that the average network distance between input neurons and their corresponding activation areas is shorter.
-Notably, applying unilateral stimulus to input neurons elicit bilateral responses, mirroring bilateral brain coordination in reality and providing additional evidence that simple network models can closely replicate reality.
-Moreover, we calculate the statistic properties and find that the network structures of two hemispheres are extremely similar.
-We also make a large spatial network visualization software.
-This research reveals network models can reach the brain’s quasi-activation pattern even with simple propagation rules.
-Besides, it provides evidence that network structure matters in brain activity pattern generation.
-Other artificial neural network research may need to make their own structure close to the real network structure.
-However, it does not perfect the model to achieve a full-activation pattern.
-Future research could enable complete simulation of brain behavior by optimizing propagation rules and automatically adjusting link weights, paving the way for achieving genuine artificial intelligence.
-6 Acknowledgement
-This work was supported by the National Natural Science Foundation of China (72025405, 72088101), the National Social Science Foundation of China (22ZDA102), the Hunan Science and Technology Plan Project (2020TP1013, 2020JJ4673, 2023JJ40685), the Innovation Team Project of Colleges in Guangdong Province (2020KCXTD040), and the Postgraduate Scientific Research Innovation Project of Hunan Province (XJCX2023095).
-The authors declare that they have no conflict of interest.
+Figure S3: Experiment of activation rate under single stimulus. We set threshold $\sigma_{0}=0.8$.
+Appendix E Impact of Perturbation Rate on Activation Areas
+The main text demonstrates that with a 10% perturbation in the network, the original realistic activation patterns vanish regardless of the model used.
+In this section, we discuss the impact of different reconnection rates on activation patterns using the degree-preserving reconnection method.
+The experiments presented in this section involve calculating patterns induced by two types of stimuli across three models under network perturbations of 1%, 3%, 5%, 10%, and 20% reconnect rates.
+We normalized each data point using the difference between the activation ratio of region $r$ and the average activation ratio $\bar{C}$ of the entire brain, that is to say, $\Delta C=C_{r}-\bar{C}$.
+The experimental results are shown in Fig.S4, where 0 represents the original, unperturbed experimental results.
+It can be observed that at the 0 point, there is a significant difference in the relative activation ratios between visual and olfactory areas under specific stimulus.
+However, this significant realistic activation patterns which we discussed in the main text disappear with a 1% network perturbation, replaced by a roughly equal activation ratio across all regions.
+This indicates that even a 1% change in network structure can lead to a rapid loss of network functionality specificity.
+Additionally, it is noted that small-scale network perturbations (1%-5%) have a lesser impact on olfactory stimulus compared to visual stimulus.
+This is because the number of neurons associated with olfactory functions is significantly fewer than those related to visual functions, making it less likely for small-scale network structure changes to affect these neurons, thereby leading to a smaller impact on functionality.
+At the same time, various horizontal comparisons indicate that while there are no significant differences among different models in forming activation patterns, more complex neuron models have a stronger ability to maintain their original activation patterns under network disturbances.
+This is because the signals generated by complex neuron models have greater redundancy and functional diversity.
+For instance, the LIF (Leaky Integrate-and-Fire) model has a voltage decay mechanism, where the voltage gradually decays in the absence of a signal.
+Even if some input signals are erroneous, as long as the proportion of errors is not very large, the decay mechanism can filter out the impact of these erroneous signals.
+These results indicate that the formation of realistic activation patterns is highly sensitive to changes in network structure, with even minor perturbations capable of disrupting this network configuration.
+This indirectly substantiates that the authentic network structure is crucial for the formation of genuine activation patterns.
+Additionally, it highlights that complex functions (such as vision) are more susceptible to network disturbances.
+(a) threshold model, visual stimulus
+(b) sigmoid model, visual stimulus
+(c) LIF model, visual stimulus
+(d) threshold model, olfactory stimulus
+(e) sigmoid model, olfactory stimulus
+(f) LIF model, olfactory stimulus
+Figure S4: Relationships of activation rate with reconnect rate. We conduct 5 repeated experiments and take the average and standard deviation.
+Appendix F Large Spatial Network Visualization
+The visualization software of experiment observation on large spatial networks is important for researchers[57][58].
+However, there is still a lack of visualization system that can observe information propagation on large spatial networks in real-time, which poses difficulties in experiment.
+To observe directly the communication progress on this super large Drosophila connectome, we develop a large spatial network visualization software.
+This software will be made available to other researchers with the author’s permission.
+The drawing part mainly employs the Canvas provided by HTML5 and we use the Three.js based on WebGL technology to achieve spatial network visualization effects, while the user interface display is constructed by Vue.js, providing users with an intuitive and interactive interface to display and analyze data.
+F.1 Front End Technical Details
+We use the Vue.js framework to build the user interface.
+Vue.js is a popular JavaScript framework that simplifies the development of web applications through componentization, enabling the creation of front-end interfaces with rich interactivity and good user experience.
+The rendering of three-dimensional complex networks relies heavily on the Three.js library.
+Three.js is a powerful JavaScript library for creating and displaying 3D graphics.
+It offers a rich set of 3D objects, materials, lights, and animations, allowing for the easy creation of high-quality 3D scenes and objects.
+To improve rendering performance, GPU acceleration technology is utilized.
+GPU acceleration takes advantage of the computational power of the Graphics Processing, significantly enhancing rendering speed and efficiency through parallel processing and stream processing modes.
+The GPU acceleration capabilities of Three.js, especially through the creation of point objects using Float32BufferAttribute, which contains their position information in three-dimensional space, greatly enhance the rendering speed and smoothness.
+F.2 Back End Technical Details
+We use the Spring Boot framework, an open-source framework based on Java for quickly building independent, production-grade Spring applications.
+With Spring Boot, backend services can be rapidly set up to provide RESTful API interfaces for frontend consumption.
+The backend also utilizes the PostgreSQL database to store network data.
+Besides, We use PostgreSQL, which is a powerful open-source relational database management system that offers a rich set of data types and robust querying capabilities.
+To store 3D network data, we leverage PostgreSQL’s JSONB type, which allows storing JSON-formatted data within PostgreSQL and provides a series of functions and operators for querying and manipulating this data.
+With the JSONB type, information about nodes and edges in the network can be stored, enabling fast querying and updating operations.
+F.3 Function Display
+Fig.SS5 is an example of the 3D visualization system.
+(a) Side view 1
+(b) Side view 2
+(c) Side view 3
+(d) Top view 1
+(e) Top view 2
+(f) Top view 3
+(g) Front view 1
+(h) Front view 2
+(i) Front view 3
+Figure S5: Visualization of Spatial Drosophila Brain Network
+All nodes’ colors can be adjusted by visible selection or input RGB code (Fig.SS6).
+Figure S6: Display of color selection
+Besides, from square to round, the shape of nodes can be adjusted even real neuron (Fig.SS7).
+Figure S7: Display of color selection
+It also can present different single areas (Fig.SS8).
+Figure S8: Display of color selection

### v4

@@
-Network Structure Trumps Neuron Dynamics: Insights from Drosophila Connectome Simulations
+Network Structure Governs Drosophila Brain Functionality
 Abstract
-Despite the success of artificial neural networks, the necessity of real network structures in simulating intelligence remains unclear.
-Utilizing the largest adult Drosophila connectome data set, we constructed a large-scale network communication model framework based on simple neuronal activation mechanisms to simulate the activation behavior observed in the connectome.
-The results demonstrate that even with simple propagation rules, models based on real neural network structures can generate activation patterns similar to those in the actual brain.
-Importantly, we found that using different neuronal dynamics models, all produced similar activation patterns.
-This consistency across different models emphasizes the crucial role of network topology in neural information processing, challenging views that rely solely on neuron count or complex individual neuron dynamics.
-Moreover, we test the influence of network reconnect rate and find that at least 1%’s reconnect rate will ruin the activation patterns appeared before.
-By comparing network distances and spatial distances, we found that network distance better explains the information propagation patterns between neurons, highlighting the importance of topological structure in neural information processing.
-To facilitate these studies, we developed real-time 3D large spatial network visualization software, bridging a crucial gap in existing tools.
-Our findings underscore the importance of network structure in neural activation and provide new insights into the fundamental principles governing brain functionality.
+How intelligence emerges from living beings has been a fundamental question in neuroscience.
+However, it remains largely unanswered due to the complex neuronal dynamics and intricate connections between neurons in real neural systems.
+To address this challenge, we leveraged the largest available adult Drosophila connectome data set, and constructed a comprehensive computational framework based on simplified neuronal activation mechanisms to simulate the observed activation behavior within the connectome.
+The results revealed that even with rudimentary neuronal activation mechanisms, models grounded in real neural network structures can generate activation patterns strikingly similar to those observed in the actual brain.
+A significant discovery was the consistency of activation patterns across various neuronal dynamic models.
+This consistency, achieved with the same network structure, underscores the pivotal role of network topology in neural information processing.
+These results challenge the prevailing view that solely relies on neuron count or complex individual neuron dynamics.
+Further analysis demonstrated a near-complete separation of the visual and olfactory systems at the network level.
+Moreover, we found that the network distance, rather than spatial distance, is the primary determinant of activation patterns.
+Additionally, our experiments revealed that a reconnect rate of at least 1‰ was sufficient to disrupt the previously observed activation patterns.
+We also observed synergistic effects between the brain hemispheres: Even with unilateral input stimuli, visual-related neurons in both hemispheres were activated, highlighting the importance of interhemispheric communication.
+These findings emphasize the crucial role of network structure in neural activation and offer novel insights into the fundamental principles governing brain functionality.
 Keywords: brain network, network communication model, drosophila connectome, activation pattern
 1 Introduction
-The neural systems of intelligent beings maintain constant communication[1, 2].
-The intricate mechanism of this communication unfolds as follows: specific neurons, often input neurons, receive stimulus signals from the external environment or other parts of the body, changing from a resting to an activated state.
-Subsequently, they relay these stimulus signals to the next tier of neurons through synapses.
+The neural systems of intelligent beings maintain continuously dynamical communication[1].
 This process gives rise to macroscopic emergence across the entire neural system.
-Such communication fosters the emergence of intelligence in all intelligent life forms, including human beings.
+Such communication fosters the emergence of intelligence in all intelligent life forms.
 Consequently, comprehending and replicating this process has become an indispensable avenue for unraveling biological intelligence and realizing artificial intelligence[2].
-Neural systems are interconnected networks[4], with neurons engaging in constant communication via synapses within the network’s topology.
-Some scientists think the brain’s unique and complex network structure is pivotal to the emergence of intelligence[5].
-Many scientists posit that network science is crucial for deciphering the mechanisms underlying the flexible regulation of neural communication[6, 7, 8].
-Determining how to leverage network science methodologies stands as a paramount challenge in contemporary neuroscience.
-However, whether this network structure determines the emergence of intelligence remains a question.
-In recent years, artificial neural networks (ANNs) have taken rapid development in enormous fields[9, 10, 11].
-A notable recent achievement is large language model (LLM).
-However, the structure of any ANNs does not follow the actual brain network structure strictly, which poses a challenge to the viewpoint that network structure inherently contains intelligence.
-Simultaneously, many scientists in machine learning adhere to the scaling law: the more parameters the network has, the greater the degree of emergent intelligence[12].
-This viewpoint suggests that intelligence is solely related to the number of parameters a model has, rather than the unique network structure between neurons.
-Under this perspective, the reason biological brains generate intelligence is due to the number of neurons, and the unique network structure is a coincidence of evolution.
-To demonstrate the importance of brain network structure, network scientists are constantly exploring and experimenting.
-An approach is brain network communication model[13].
-Scientists wants to use brain network communication model to understand and describe the interaction and communication between different parts of the brain.
-This model aids in studying the structural and functional aspects of neuronal networks and how they coordinate to accomplish complex cognitive processes[14].
-The first challenge of this approach is obtaining a complete and accurate mapping of the brain’s structure.
-For example, current imaging technologies, such as MRI, fMRI, and electron microscopy, have limitations in resolution and speed.
-While MRI and fMRI provide non-invasive imaging, their spatial resolution is not sufficient to map individual neurons and synapses.
-On the other hand, techniques like electron microscopy, which can achieve higher resolution, are extremely slow and generate enormous amounts of data that are challenging to process[15].
-Modeling from the micro scale is necessary because modeling at the level of brain regions or nerve fibers does not include the most basic structural information and lacks the most complete physical foundation.
-Moreover, current models often rely on prevalent metrics such as signaling cost and activation time to assess model quality[16].
-Yet, these metrics fall short of verifying whether a network model accurately reflects the brain’s actual activation patterns, which is the fundamental goal of brain simulation.
-There is a need for evaluation criteria that can affirm whether network communication models genuinely mirror the brain’s communication processes.
-Based on facts and challenges above, benefiting from a largest-to-date adult Drosophila connectome data[17], this study attempts to investigate how the type of structure of neurons could affect the activation patterns of a brain.
-Our approach involves crafting a large scale network communication model framework based on simple mechanisms, designed to simulate the activation behavior observed in the connectome.
-Specifically, we establish the network representation of the connectome.
-Then we design a large scale threshold propagation model of nerve cells: when a neuron received enough stimuli, it activates and subsequently transmits signals to other neurons.
-After a few iterations, the whole network emerges an activation pattern.
-We use threshold model and sigmoid model as the basic neuron model.
-We also design an improved LIF model as the basic neuron model for comparison.
-Additionally, we propose an evaluation criteria based on the ratio of actually activated neurons to the expected number of activation.
-Experiments reveals that, from threshold , sigmoid, and LIF neuron model, no matter activation mechanism is, our model has significant activation in the type of neurons that should be affected by the stimulus and slight activation in irrelevant neurons.
-Moreover, to supply more evidence of the vitalness of network structure, we disturb the network by the way of degree-preserving rewiring and and find the disturbed network can’t generate any correct activation pattern.
-In addition, we test the sensitivity of disturb degree.
-For details, we record the activation rate of areas under different reconnect rates and find that 1%’s random change of network structure will cause a complete change in function.
-To prove the network structure, rather than the spatial structure, matters in activation pattern generation, we calculate the average network and spatial distance of input and activation neurons and results reveal that the average network distance between input neurons and their corresponding activation areas is shorter than unrelated areas, which is a strong evidence of network matters in message propagation in brain.
-Furthermore, when providing unilateral stimuli to input neurons, we observe a bilateral response in the Drosophila connectome, aligning with reality.
-Then we calculate the statistics properties and find that the left and right hemispheres have extremely similar degree, clustering coefficient, and eigenvector centrality vector coefficient.
-The vectors’ Pearson coefficient are up to 0.9986, 0.9988, and 0.9991, respectively.
-We also developed a 3D network visualization software see in appendix F, utilizing HTML5’s Canvas and Three.js based on WebGL technology for large spatial network visualization effects.
-This tool enables real-time observation of each neuron’s communication process.
+Neural systems are often modeled as interconnected networks[3], with neurons engaging in constant communication via synapses within the network’s topology.
+To elucidate the relationship between the network itself and its dynamics in a neural network, various models have been developed, spanning from the cellular level to the whole brain.
+For example, neuronal dynamic models like Leaky Integrate-and-Fire (LIF) model[4] are designed to describe the electrical activity and dynamic behavior of neurons[5].
+These models are characterized by the following features: they typically use a set of differential equations to describe the electrophysiological behavior of neurons and usually contain a set of parameters that need to be measured experimentally or estimated.
+On a larger scale, a vast array of isolated neuronal dynamics, ranging from several to millions, are interconnected and parameter-optimized to simulate the local or global dynamics of the brain[6].
+Due to the requirements of commonly used algorithms like backpropagation on network structures, these methods typically do not fully adhere to the structures of real neural networks.
+Another category of network-based methods applies propagation models on complex networks to neural networks, thereby describing information propagation and emergent mechanisms within neural networks.
+These methods are collectively referred to as brain network communication models.
+Such approaches typically involve substantial simplification of neuron models, resulting in fewer parameters, but also leading to the loss of some electrophysiological details of neurons[7].
+These two seemingly opposing approaches illustrate that it remains unresolved whether the activation mechanisms of neurons or the connectivity between neurons play the more dominant role.
+This question is fundamental to understanding how intelligence emerges from neural systems.
+Additionally, in recent years, as an innovative way to modify intelligence, artificial neural networks (ANNs) have achieved significant advancements in numerous fields.
+[8].
+Despite these achievements, the structure of ANNs does not strictly mimic the actual neural network structure of biological brains.
+This discrepancy raises questions about the hypothesis that the inherent structure of a network is crucial for intelligence.
+Furthermore, the scaling law in the field of machine learning posits that the more parameters an ANN has, the greater the degree of emergent intelligence it can exhibit[9].
+These theory hints that the quantity of neurons (parameters) might be a more significant factor in the development of intelligence than the precise arrangement of these neurons in a network, suggesting that the specific network structures found in biological systems may not be essential for the emergence of intelligence.
+The observation of neural network structures is crucial for interpreting intelligence from a network perspective.
+Current neuroimaging techniques, such as Magnetic Resonance Imaging (MRI) and functional MRI (fMRI), offer non-invasive exploration but lack the spatial resolution necessary to resolve individual neurons and synapses.
+However, these measurement methods face challenges in obtaining information about the positions and connectivity of neurons in the nervous system, making it difficult to understand the nervous system at the neuron-synapse level[10].
+Employing electron microscopy (EM) and deep learning techniques for automated neuron reconstruction, brain atlases of a range of animals, including Drosophila larvae[11], have been mapped.
+A recent study[12] published an adult Drosophila connectome data set, comprising nearly 130,000 neurons and 50,000,000 synapses, with details of neuronal network structure at unforeseen resolution.
+With full proofreading of the central brain and approximately 80% completion of the optic lobesis presented, the data represents the largest whole-brain connectome of an adult Drosophila to date and provides an unprecedented opportunity for comprehensive network analysis and topological characterization of the Drosophila brain.
+Building on the outlined challenges and opportunities, this study utilized the aforementioned Drosophila connectome data set[12] to examine the importance of neuronal network’s structure in simulating brain activation patterns.
+We constructed a large-scale network communication model framework based on simple mechanisms to simulate activation behavior observed in the connectome.
+Initially, we established a network representation of the connectome.
+We then developed a large-scale network communication model where neurons activate upon receiving sufficient stimuli and transmit signals to other neurons.
+This iterative process generated a global activation pattern.
+Our analysis employed threshold and sigmoid models as baselines, with a LIF model for comparison.
+We proposed an evaluation criterion based on the ratio of actual to expected neuron activations.
+Additionally, we developed a 3D network visualization tool (see Appendix F), using HTML5’s Canvas and Three.js based on WebGL technology.
+This software enables real-time observation of each neuron’s communication process in three dimensional space.
 2 Results
-First, we report the basic statistics of Drosophila connectome.
-After that, to test the performance of our model in terms of realism on the data set and determine whether different artificial mechanisms can achieve similar activation patterns when the network structure changes, we give different stimuli and record the activation pattern of different neurons.
-In addition, we give a unilateral stimulus (stimulus one hemisphere) and record the bilateral response of neurons.
-Besides, to add more neurons’ real dynamic details, we also add an improved LIF model as the basic neuron dynamic model and record the activation patterns under stimuli.
-Finally, to prove the network structure matters in brain network communication, we calculate the average physical and network distance between different areas.
-2.1 Experimental Settings
-The basic methodology of the experiment is described as follows.
-In this study, a continuous stimulus signal (binary 0-1 signal) is applied to a subset of input neurons (such as visual or olfactory neurons) in the brain.
-Subsequently, we compute the signal’s propagation process within the brain and record the responses of both intermediate and output neurons.
-The evaluation criterion is determined by calculating the ratio of neurons that ought to be activated to those that are actually activated in a specific area (see as Equ.26 in appendix B).
-Given that the brain functions as a type of shallow neural network[53], we document the outcomes at an iteration step size of 5.
-Moreover, as the input remains constant and no randomization parameters have been introduced, the outcomes of each experiment are consistent, eliminating the need for repetition.
-For messages in related neurons, please read appendix C.
-In linear threshold model, we set $\sigma_{0}=0.8$.
-In the simulation of the LIF model, numerous hyperparameters need to be used.
-All the hyperparameters used are listed in the following table.
-Biological hyperparameters are set according to real biologic experiment test[54].
-Current proportion const is the optimal parameter obtained under equidistant search.
-Table 1: List of Hyperparameters
-Hyperparameter Description Value
-$\tau$ Membrane time constant 15.0 ms
-$V_{\text{rest}}$ Resting membrane potential -65 mV
-$V_{\text{th}}$ Spike threshold -45 mV
-[…]
-$dt$ time step 0.55 ms
-$\alpha$ Current proportion const 0.03/0.045 (visual/olfactory)
-2.2 Visualization and Statistic Properties of Drosophila Connectome
-In this section we first present the visualization of Drosophila Connectome (Fig.1(a)).
-Other visualization cases and methods of visualization are presented in appendix F.
-Visually, the shape of the brain appears to be physically similar, but whether the network structures are similar remains unknown.
-So we further investigate the network statistics of two hemispheres.
-We calculated the degree, clustering coefficient, and eigenvector centrality distribution for every node in each hemisphere (Fig.1(d), (e) and (f)).
-The Pearson correlation coefficient of degree, clustering coefficient, and eigenvector centrality vector in each hemisphere is 0.9986, 0.9988, and 0.9991, respectively.
-According to Fig.1(d), (e) and (f), these statistical indicators for both hemispheres are extremely similar, indicating that the brain exhibits strong symmetry in network structure.
-From Fig.1(c) we can see that the degree, i.e., the number of connections to the nodes in the network, exhibits a long-tail distribution, implying that while a large number of nodes possess few connections, a small number of nodes contains a large number of connections[37].
-It implies the real neural network have some same statistic properties as social, electrical and transportation networks.
-Also, we plot the correlation curves between degree and cell surface area, volume, and cell length (Fig.1).
-The results indicate that the degree distribution is logarithmically linearly correlated with these characteristics.
+2.1 Structural Properties of Drosophila Connectome
+The structure of the Drosophila connectome is presented in Fig.
+1(a).
+Visually, the shape of most of the functional regions of the brain appears to be physically similar.
+From Fig.
+1(b) we can see that the degree, i.e., the number of connections to the nodes in the network, exhibits a long-tail distribution, implying that while a large number of neurons possess few synapses, a small number of neurons contains a large number of synapses.
+For example, more than 20% of the nodes have a degree greater than 50, while only 0.3% of the nodes have a degree greater than 500.
+This is consistent with previous research on functional brain networks[13].
+To investigate whether the two hemispheres’ network structures are similar in network statistic properties, we conducted a comprehensive network analysis on the two hemispheres.
+We calculated the degree, clustering coefficient, and eigenvector centrality distribution for every node in each hemisphere.
+According to Fig.
+1(c)-(e), these statistical indicators for both hemispheres are extremely similar, indicating that the brain exhibits strong symmetry in network structure.
+Also, we plotted the correlation curves between degree and cell length, cell surface area, and cell size (Fig.
+1(f)).
+The significance coefficients $R$ are 0.85 ($p=0.00$), 0.86 ($p=0.00$), 0.80 ($p=0.00$), respectively.
+The results indicate that the degree is logarithmically linearly correlated with these characteristics.
 This logarithmic relationship reveals a fundamental principle in the organization of neural networks in the brain.
-To elaborate, as the surface area, volume, or length of a neuron increases, the number of synapses (represented by the degree) it forms grows logarithmically rather than linearly.
-This non-linear scaling is crucial for maintaining efficient information processing in the brain while preventing an explosion in connectivity that would be computationally and energetically costly.
-These characteristics of the brain can explain Weber-Fechner’s law[47]: for any sensory pattern, the perceived intensity is a logarithmic function of the physical intensity.
+Figure 1: Visualization and statistic properties of Drosophila connectome data set.
+(a) is a visualization of Drosophila connectome.
+The different colors represent different functional regions.
+(b) shows the degree distribution of the network.
+(c), (d) and (e) plot the distribution of degree, clustering coefficient, and eigenvector centrality in each hemisphere, respectively.
+(f) represents the correlation between degree distribution and neuron characteristics.
+We use nodes to represent neurons’ cell bodies and use edges to represent the synapses when calculated statistic properties.
+2.2 Activation Patterns under Different Type of Stimuli
+In this section, we investigate the impact of different activation mechanisms on network activation patterns by utilizing various network activation mechanisms and observing the changes in activation rates across different areas of the network.
+Simultaneously, we examine the effect of global perturbations in network structure on activation patterns by altering the network architecture.
+The optic and visual_projection area are related to vision.
+ALLN, ALIN, ALON and ALPN are related to olfaction.
+Other supplemental messages of related neurons are presented in Appendix B.
+Fig.
+2 illustrates the activation patterns of the large-scale network communication model (LSNC) using threshold, sigmoid, and LIF activation mechanisms.
+We also evaluated the model’s performance after introducing perturbations.
+For a detailed description of these models, please refer to Appendix A.2.
+From left subgraphs of Fig.
+2 (a) and (b), it can be observed that the network model generates quasi-real activation patterns under specific stimuli: It is evident that visual (olfactory) stimuli result the most significant activation on visual (olfactory) neurons such as optic and visual projection neurons (ALLN, ALIN, ALPN, ALON neurons) while activating other neuron types minimally.
+Our findings indicate that different activation mechanisms can yield comparable activation patterns.
+In response to visual stimuli, all three models displayed notable activation in visual-related areas, whereas other areas exhibited minimal activation (left subgraphs of Fig.
+2 (a), (c), and (e)).
+The optic area, for instance, recorded activation rates of 36.39%, 4.69%, and 9.69%.
+Under olfactory stimuli, the network also exhibits similar activation patterns in olfactory-related areas, with average activation rates of 94.07%, 78.01%, and 69.75%, respectively (left subgraphs of Fig.
+2 (b), (d) and (f)).
+This suggests that the connectivity between neurons is more critical than the activation mechanisms of individual neurons.
+However, in the perturbed models (right subgraphs of Fig.
+2 (a)-(f)), the original quasi-real activation patterns disappear.
+No significant differences in activation regions were observed across all activation modes and types of stimuli.
+For instance, in the model with threshold activation mechanism, the maximum difference of all areas in the activation ratios under visual stimuli was only 3.162%, and the range of regional activation ratios versus the average activation ratio was only 6.32% (right subgraph of Fig.
+2(a)), which indicates that the heterogeneity between different regions completely disappears.
+Additionally, in the sigmoid and LIF models, no significant activation was observed in any area.
+Despite variations in activation rates among the three models, the substantial disparity between the activation ratios of relevant and irrelevant areas suggests the emergence of comparable quasi-real activation patterns across all models.
+(a) threshold, visual
+(b) threshold, olfactory
+(c) sigmoid, visual
+(d) sigmoid, olfactory
+(e) LIF, visual
+(f) LIF, olfactory
+Figure 2: Comparative test of intrinsic neurons’ activation.
+For disturbing model, we conduct 5 repeated experiments and take the average and standard deviation.
+The dotted lines represent average activate rate in brain.
+In disturbing model, we set reconnect rate $p=20\%$ and $\sigma$ = 0.8.
+2.3 Separation of Visual and Olfactory Related Areas
+To further investigate the causes of the above observed dynamic pattern that is independent of neuron models, we calculated the synapse count (number of connections) between various regions of the visual and olfactory systems and visualized the results as a heat map (Fig.
+3(a)).
+The findings reveal a significant segregation between the visual and olfactory areas.
+Specifically, within the visual and olfactory areas, there are 13,309,161 and 882,170 synapses, respectively.
+However, only 1,837 synapses exist between these two areas.
+These interconnecting synapses constitute only approximately 0.0138% and 0.208% of the total synapses within the visual and olfactory areas, respectively.
+Fig.
+3(b) represents the schematic diagram of the connection relationships in the neuronal visual and olfactory systems.
+We use nodes to represent functional areas and use edges to represent synapses between areas.
+The thickness of edges represents the number of connections.
+This diagram indicates that the number of connections within regions is significantly greater than the number of inter-regional connections.
+These indicate that the network structure between these two functional systems is almost entirely separate.
+Figure 3: Heatmap and schematic diagram illustrating the connections between areas related to vision and olfaction.
+(a) presents the heatmap, while (b) shows the schematic diagram of the connection relationships.
+2.4 Distances between Input and Activation Areas
+To further investigate whether this separation phenomenon is caused by greater spatial distance, we calculated the average network distance and spatial distance between different areas of the brain network.
+The results are presented in Tab.
+1.
+The analysis of Tab.
+1 showcases that spatial distances among neurons do not straightforwardly translate to ease of activation.
+While visual input neurons across all intrinsic areas exhibit similar average spatial distances, quasi-real activation patterns manifest predominantly in specific neural types (optic and visual_projection).
+Conversely, the average network distances between input neurons and target neurons is notably smaller compared to irrelevant neurons.
+Specifically, visual input neurons exhibit network distances to optic and visual_projection neurons as 5.7018 and 5.7432, respectively, while distances to neurons irrelevant to olfactory (ALLN, ALIN, ALON, and ALPN) are measured at 6.6442, 6.2350, 6.1506, and 6.5380, respectively.
+This demonstrates that neurons corresponding to visual inputs maintain a generally closer network distance relative to unrelated neurons.
+Furthermore, neurons associated with olfactory inputs maintain network distances approximately half that of unrelated neurons.
+These observations suggest that the network distance between input neurons and their specific activation areas is more significant than spatial distance, underscoring the importance of network structure over spatial structure in defining the quasi-real activation patterns of brain networks.
+Table 1: Spatial and network distance between input areas and functional areas.
+Input type Distance type optic visual_projection ALLN ALON ALPN ALIN
+Visual input Spatial distance (nm) $365012$ $354725$ $351919$ $346925$ $348517$ $348901$
+Network distance (edge) $\mathbf{5.7018}$ $\mathbf{5.7432}$ $6.6442$ $6.1506$ $6.5380$ $6.2350$
+Olfactory input Spatial distance (nm) $302690$ $251144$ $110827$ $123179$ $115089$ $173345$
+Network distance (edge) $5.7967$ $5.1706$ $\mathbf{3.4458}$ $\mathbf{3.6312}$ $\mathbf{3.6820}$ $\mathbf{3.3845}$
+3 Discussion
+Our study, which utilized a large-scale network communication model derived from the Drosophila connectome, highlights the crucial importance of neural network architecture in generating brain-like activation patterns.
+We first analyzed the structure of the Drosophila brain network, including the basic statistical properties and the symmetry of the network.
+An interesting phenomenon observed is the correlation between neuronal degree and geometric size: Fig.
+1(f) demonstrates that the number of synapses formed by a neuron (represented by degree) increases logarithmically, rather than linearly, with respect to its surface area, volume, or length.
+These characteristics of the brain can be explained by the Weber-Fechner’s law: for any sensory pattern, the perceived intensity is a logarithmic function of the physical intensity.
 Specifically, the size, length, and area changes can only cause a logarithmic change in the number of synapses the neuron has.
-This means that as a stimulus becomes stronger, the increase in neural activity or connectivity is proportionally smaller, mirroring the logarithmic perception of intensity described by Weber-Fechner’s law.
-(a) Visualization of Drosophila Connectome
-(b) Correlation between degree distribution and neuron characteristics
-Figure 1: Visualization and statistic properties of Drosophila connectome data set
-These statistic results are a valuable supplement to previous findings[48] and are crucial for for improved understanding of the Drosophila brain connectcome structure.
-2.3 Activation Patterns under Different Type of Stimuli
-Primarily, to validate the efficacy of our model in signal transmission between input and output neurons, we showcase the activation patterns of intermediate and output neurons throughout the information transmission process.
-This demonstration serves to substantiate that our model accurately reflects real-world scenarios, shaping a quasi-real activation pattern.
-The specific idea for the experiment is following: we stimulate different types of neurons (such as visual input neurons, olfactory neurons, etc.), and record the rate that neurons that are activated in different brain areas.
-If a specific stimulus causes a significantly greater response than the average in the brain area that it is theoretically supposed to activate, then it is considered that the model can generate a quasi-real activation pattern in the brain.
-We give 2 types of input neurons (visual, olfactory) stimulus and record the activation patterns of 6 type of neurons: optic (visual), visual_projection (visual), ALLN (olfactory), ALIN (olfactory), ALPN (olfactory) and ALON (olfactory).
-When we give neuron a stimulus, we set the state of neuron as 1.
-To demonstrate the importance of network structure, we add a comparative model that randomly reconnects networks.
-Specifically, we perform degree-preserving rewiring on a certain proportion of edges in the brain, which means perturbing the network structure while ensuring that the degree of each node remains unchanged.
-In Fig.2 we set reconnect rate $p=10\%$.
-For its sensitivity analysis, please read appendix S4.
-Fig.2 shows the activation patterns of different models.
-From Fig.
-2(a), (e), and (i), it shows visual stimulus cause a strong impact (significantly more than average) on visual neurons (optic, visual projection) while causing a slight impact (significantly less than average) on other type of neurons under all three kinds of models.
-Specifically, under visual stimulus, in response to visual stimuli, the optic (vision related area) exhibit a 36.39%, 4.69% and 9.69% ’s activate rates, respectively.
-At the mean time, no olfactory-related neurons are activated.
-In addition, olfactory stimulus cause a strong impact (significantly bigger than average level) on olfactory neurons (ALLN, ALIN, ALPN, ALON), reaching the activation rate by 94.07%, 78.01%, and 69.75%, respectively.
-And it cause a slight impact (significantly less than average) on other type of neurons.
-For example, it only activates 0.04% of optic and 0.79% of visual_projection in threshold model (Fig.2(c)).
-Although the activation rates of the three models in specific areas differ (due to different activation mechanisms), the significant difference in the activation ratio between correlated and uncorrelated regions indicates the emergence of similar quasi-real activation patterns across all models.
-This suggests that different activation mechanisms are not the fundamental reason for changes in network functionality.
-As for the disturbing models, in Fig.2 (b), (d), (f), (h), (j) and (l), the quasi-real activation patterns totally disappear.
-Across all activation modes and all types of stimuli, no significant differences in activation regions were observed.
-For example, in the threshold model, the maximum difference in the activation ratios of all areas under visual stimuli was only 3.162%.
-Moreover, the range of the regional activation ratios versus the average activation ratio was only 6.32%.
-Additionally, in the sigmoid and LIF models, no significant activation was observed in any area, indicating that stimuli under network disturbance did not lead to any significant response in any area.
-These results indicate that no matter neuron model is, it will lead to similar activation patterns which correspond to reality.
-After disturbing the network structure, the original quasi-real activation patterns in all models completely disappear.
-It means the real network structure matters than activation mechanisms in generating the quasi-real activation patterns.
-Different activation mechanisms are not the primary cause of this realistic activation mechanism; it is the brain’s unique network structure that plays the crucial role.
-This aligns with the studies of complex systems: individual capabilities are not the determining factors in complex systems; what matters is the organization and interaction structure among individuals.
-This perspective highlights the importance of individual relationships and the overall system structure in determining emergent phenomena.
-(a) threhold, visual
-(b) threhold, visual (after disturb)
-(c) threhold, olfactory
-(d) threhold, olfactory (after disturb)
-(e) sigmoid, visual
-(f) sigmoid, visual (after disturb)
-(g) sigmoid, olfactory
-(h) sigmoid, olfactory (after disturb)
-(i) LIF, visual
-(j) LIF, visual (after disturb)
-(k) LIF, olfactory
-(l) LIF, olfactory (after disturb)
-Figure 2: Experiment of intrinsic neurons’ activation. For disturbing model, we conduct 5 repeated experiments and take the average and standard deviation.
-2.4 Distances between Input and Activation Areas
-To further investigate how the network structure, rather than the spatial distribution of neurons, or the parameters of the network model or activation mechanisms, predominantly determines the dynamics of brain activity, we calculated the average network distance and spatial distance between different areas of the brain network.
-The optic and visual_projection area are related to vision and ALLN, ALIN, ALON, ALPN, LHCENT and LHLN are related to olfactory.
-The results are presented in Tab.
-2.
-From Tab.
-2, it is evident that neurons which are closer in spatial distance are not necessarily easier to activate.
-For example, all areas and visual input neurons have similar average spatial distances, but the quasi-real activation patterns can only be observed in specific neural types (optic, visual_projection).
-In contrast, the average network distance between input neurons and neurons should be activated is closer than irrelevant neurons.
-For instance, the distance between visual input neurons and optic, visual_projection are 5.7018 and 5.7432, respectively.
-And the distance between visual input neurons and ALLN, ALIN, ALON, ALPN, LHCENT, LHLN which are irrelevant with vision are 6.6442, 6.2350, 6.1506, 6.5380, 6.3094 and 6.8206, respectively.
-The network distance of neurons related to visual input neurons is generally closer than the distances of unrelated neurons.
-Besides, The average network distance of neurons related to olfactory input neurons is approximately twice as close as the distance of unrelated neurons.
-These results indicates that compared to spatial distance, the average network distance between input neurons and their corresponding activation areas is closer, meaning that compared to spatial structure, network structure is more important, determining these quasi-real activation patterns of the brain network.
-This results reflect the importance of network structure in neural networks, that is, the ability to reproduce certain functions of the brain by considering the network structure only.
-Table 2: Spatial and network distance between input and main areas. The unit of spatial distance is nm and the unit of network distance is the number of edges in the shortest path.
-distance type Neuron type Area Name optic visual_projection ALLN ALIN
-Spatial Distance Visual Input Neurons $365012$ $354725$ $351919$ $348901$
-Olfactory Input Neurons $302690$ $251144$ $110827$ $173345$
-ALON ALPN LHCENT LHLN
-Visual Input Neurons $346925$ $348517$ $362825$ $366013$
-Olfactory Input Neurons $123179$ $115089$ $212515$ $246719$
-optic visual_projection ALLN ALIN
-Network Distance Visual Input Neurons $5.7018$ $5.7432$ $6.6442$ $6.2350$
-Olfactory Input Neurons $5.7967$ $5.1706$ $3.4458$ $3.3845$
-ALON ALPN LHCENT LHLN
-Visual Input Neurons $6.1506$ $6.5380$ $6.3094$ $6.8206$
-Olfactory Input Neurons $3.6312$ $3.6820$ $3.9557$ $3.9758$
-3 Discussion
-This study, through the construction of a large-scale network communication model based on the Drosophila connectome, reveals the critical role of neural network structure in generating brain-like activation patterns.
-Our findings suggest that models based on comprehensive neuron network data can approach the brain’s quasi-real activation patterns even with simple artificial rules.
-This paves the way for understanding biological intelligence, indicating that the structure of biological neural systems may offer valuable reference for designing efficient, fast-adapting intelligent systems.
-Our research emphasizes the importance of network structure, compared to spatial structure or neuronal activation mechanisms, in determining brain activity patterns.
-This finding has significant implications for artificial neural network research, suggesting that future AI designs may need to more closely mimic real brain network structures.
-Furthermore, the large spatial network visualization software we developed provides a new tool for neural network research, facilitating intuitive observation and analysis of complex neural network dynamics.
+These results suggest that although synaptic plasticity lead to a continuous increase in the number of synapses between neurons and other cells, this logarithmic relationship makes it difficult for the number of synapses formed by a neuron to increase indefinitely.
+This may account for the overall power-law distribution observed in Fig.
+1(f).
+To examine the impact of neuronal activation mechanisms on the generation of realistic network patterns, we simulated three activation mechanisms for specific inputs.
+Fig.
+2 suggests that, the real network structure is more critical than activation mechanisms in generating quasi-real activation patterns.
+The brain’s unique network structure plays a crucial role in this realistic activation mechanism, aligning with studies on complex systems where the organization and interaction structure among individuals, rather than individual capabilities, determine emergent phenomena[14].
+To elucidate the influence of network structure on function, we analyzed the topology of the network (Fig.
+3 and Tab.
+1).
+These results collectively suggest that network architecture, rather than spatial structure or activation mechanisms, is the primary determinant of the Drosophila connectome’s ability to generate quasi-real activation patterns, owing to its unique connectivity profile.
+To test the impact of network perturbation ratio on its functionality, we calculated activation patterns with different reconnect rates (Appendix S5).
+These results indicate that the formation of realistic activation patterns is highly sensitive to changes in network structure, with even minor perturbations capable of disrupting this network configuration.
+This indirectly substantiates that the authentic network structure is crucial for the formation of genuine activation patterns.
+Additionally, it highlights that complex functions (such as vision) are more susceptible to network disturbances.
+Conversely, Fig.
+S4 shows that the network retains its original performance under a region-based attack.
+One explanation for these findings is that random rewiring thoroughly alters the network topology, potentially disrupting original functional modules and pathways, which leads to disordered information transmission paths and significantly impacts activation patterns.
+Additionally, because rewiring is global and random, it induces unpredictable perturbations that undermine redundancy and fault tolerance mechanisms, rendering the network more sensitive to such global disturbances.
+In contrast, the reason the network’s dynamic properties remain relatively stable after the deletion of nodes within a specific region is that this perturbation is localized and has a minor impact on the global topology.
+Moreover, random rewiring may affect critical nodes or key pathways within the network, leading to significant functional impairment of the overall network.
+However, regional node deletion, if involving non-critical nodes and edges, may not significantly impact the network’s functionality, allowing it to maintain normal activation patterns.
+To evaluate the impact of the observed network structural symmetry (Fig.
+1(a), (c), (d), and (e)), we input visual signals separately to the left and right hemispheres and observed the effects.
+The findings in Appendix D reflect real-world phenomena: Although the left and right hemispheres possess certain functional specializations, they also exhibit synergistic effects and typically collaborate to fulfill specific functions [15].
+For example, during complex cognitive tasks, both hemispheres coordinate their efforts to achieve task completion.
+The experimental results indicate that similar synergistic effects emerge even when only the neurons on one side of the brain are stimulated, indirectly validating the model’s effectiveness.
+Moreover, the network maintains similar synergistic effects even in the face of large-scale localized attacks, indirectly demonstrating the network’s robust resilience.
+Finally, by analyzing the connections between different regions, it is found that the cause of these synergistic effects is not the connections between visual input and contralateral visual processing areas, but rather the extensive connections within the visual processing regions themselves.
+This provides a new perspective for understanding the brain’s synergistic effects.
+These findings has significant implications for artificial neural network research, suggesting that future AI designs may need to more closely mimic real brain network structures.
 Future research directions may include optimizing propagation rules and automatically adjusting link weights to achieve complete simulation of brain behavior.
@@
 for their invaluable help in sharing the Drosophila connectome data.
-This work was supported by the National Natural Science Foundation of China (72025405, 72088101), the National Social Science Foundation of China (22ZDA102), the Hunan Science and Technology Plan Project (2020TP1013, 2020JJ4673, 2023JJ40685), the Innovation Team Project of Colleges in Guangdong Province (2020KCXTD040), and the Postgraduate Scientific Research Innovation Project of Hunan Province (XJCX2023095).
+This work was supported by the National Natural Science Foundation of China (72025405, 72088101, 72301285, 72001211), the National Social Science Foundation of China (22ZDA102), and the Natural Science Foundation of Hunan Province (2023JJ40685).
 The authors declare that they have no conflict of interest.
 APPENDIX
-Appendix A Related Works
-Various models have been developed to mimic or understand aspects of brain function, ranging from the cellular level to the whole brain.
-In this section, we review the models that aim to simulate the brain’s behavior.
-The review includes neural dynamic model that describes how action potentials in neurons are initiated and propagated.
-In addition, we introduce the large scale brain simulating methods.
-At last, we introduce the network communication models which refers to a computational or theoretical framework designed to understand and simulate how different parts of the brain communicate with each other.
-A.1 Neuron Dynamic Model
-Neuronal dynamic model is a mathematical framework designed to understand and simulate the electrical activity and dynamic behavior of neurons.
-These models aim to capture the fundamental mechanisms by which neurons process and transmit electrical signals through their membranes, often by using differential equations and computational techniques.
-In past years, neuroscientists have aimed to describing the behavior of neurons by dynamic model.
-One of the most famous works is the Hodgkin-Huxley (HH) model[18].
-It includes a set of differential equations that describe the changes in membrane potential and the dynamics of ion channels.
-The key variables in the model are the membrane potential $V$, and the gating variables $m$, $n$, $h$ that represent the probability of ion channels being open.
-The currents are given by:
-$I_{Na}$ $=\bar{g}_{Na}m^{3}h(V-E_{Na}),$ (1)
-$I_{K}$ $=\bar{g}_{K}n^{4}(V-E_{K}),$ (2)
-$I_{L}$ $=\bar{g}_{L}(V-E_{L}),$ (3)
-where $\bar{g}_{Na}$, $\bar{g}_{K}$ and $\bar{g}_{L}$ are the maximum conductances for sodium, potassium, and leak channels.
-$E_{Na}$, $E_{K}$, and $E_{L}$ are the reversal potentials for sodium, potassium, and leak channels.
-The gating variables obey first-order kinetics and are described by the following differential equations:
-$\frac{dm}{dt}$ $=\alpha_{m}(1-m)-\beta_{m}m,$ (4)
-$\frac{dn}{dt}$ $=\alpha_{n}(1-n)-\beta_{n}n,$ (5)
-$\frac{dh}{dt}$ $=\alpha_{h}(1-h)-\beta_{h}h,$ (6)
-where $m$, $n$, and $h$ are the gating variables representing the activation and inactivation states of the ion channels.
-$\alpha_{m}$, $\alpha_{n}$, $\alpha_{h}$, $\beta_{m}$, $\beta_{n}$, and $\beta_{h}$ are voltage-dependent rate constants for the gating variables.
-These rate constants are defined by empirical equations based on experimental data.
-HH model accurately describes the electrophysiological processes of neuronal activity through mathematical equations, allowing scientists to understand the fundamental mechanisms of a group of neurons’ behavior at the molecular and cellular levels.
-Another classic theory is cable theory[19].
-Cable theory is a mathematical model used to understand how electrical signals propagate along neurons, especially in their dendrites and axons.
-This theory applies the principles of electrical circuits to describe how the passive properties of the neuron influence signal transmission.
-It’s particularly useful in explaining the conduction of graded potentials, such as synaptic potentials and electrotonic (passive) spread of membrane potential changes.
-The core equation of cable theory is derived from the cable equation, which is a form of the second-order linear partial differential equation. It can be expressed as:
-$\frac{\partial^{2}V}{\partial x^{2}}=\frac{1}{r_{m}}\left(\frac{\partial V}{\partial t}+\frac{V}{r_{c}}\right),$ (7)
-where $V$ is the membrane potential, $x$ is the distance along the cable, $t$ is time, $r_{m}$ is the membrane resistance per unit length, $r_{c}$ is the cytoplasmic (axial) resistance per unit length.
-This equation describes how the voltage $V$ changes along the length of the neuron $x$ and time $t$.
-Cable theory has several important applications in neuroscience, including understanding signal decay, neural modeling and synaptic integration.
-Despite its usefulness, cable theory has limitations.
-First, real neurons have complex geometries.
-Cable theory don’t describe well of these complex geometries.
-This can lead to inaccuracies in predicting real neuronal behavior.
-Besides, cable theory primarily addresses passive electrical properties and does not account for active properties like action potentials, which are crucial for neural communication.
-Apart from that, researchers proposed many variants of neuron dynamic models[20][21].
-These models serve to elucidate the physiological activity of neurons and contribute to the description of neuron behavior.
-However, these models has some limitations: these equations often contain a multitude of parameters.
-In instances where measurement techniques fall short of quantifying each parameter individually, estimation becomes a necessity.
-Spiking Neural Networks (SNNs)[22] is designed to optimize the parameters in neural dynamic models.
-Unlike their traditional counterparts, which process information through continuous values, SNNs utilize discrete events known as spikes.
-These spikes, brief in duration, mimic the communication method of neurons within the brain, facilitating the transmission of information.
-One of the simplest and most commonly used models to describe neuron dynamics in SNNs[22] is the Leaky Integrate-and-Fire (LIF) model[23].
-The LIF model captures the essential characteristics of spike generation in neurons.
-The membrane potential $V(t)$ of a neuron at time $t$ is described by the following differential equation:
-$\tau_{m}\frac{dV(t)}{dt}=-[V(t)-V_{rest}]+RI(t),$ (8)
-where $\tau_{m}$ is the membrane time constant, which determines how fast the membrane potential decays towards the resting potential $V_{rest}$ in the absence of input.
-$V_{rest}$ is the resting membrane potential.
-$R$ is the membrane resistance.
-$I(t)$ is the input current at time $t$.
-When $V(t)$ reaches a certain threshold $V_{threshold}$, the neuron fires a spike, and $V(t)$ is reset to $V_{rest}$ or another predefined value.
-After firing, the neuron may enter a refractory period during which it cannot spike, regardless of the input.
-SNNs have enabled researchers to simulate networks comprising millions of neurons, tackling a wide array of tasks ranging from image classification and natural language processing to controlling robotic arms[24][25][26][27].
-Despite their advances, SNNs face significant challenges.
-One primary issue is that the SNN’s structure often draws inspiration from functional brain networks rather than accurately replicating the intricate neural connections observed at the microscale.
-In terms of the connection method of artificial neurons, SNNs usually use a structure similar to that of ANNs.
-Consequently, these models resort to architectures akin to those found in artificial neural networks, hindering their ability to authentically emulate brain functions.
-Additionally, SNNs are marked by their low computational efficiency, with a single forward propagation frequently requiring several seconds to complete.
-These limitations underscore the difficulties in considering SNNs as the optimal approach for brain simulation in their current state.
-As for the large scale brain simulation model aims to simulate the whole brain’s behavior, Eugene et al.[28] presents the first large scale brain simulation model based on experimental measures in several mammalian species.
-The model is founded upon the global thalamocortical anatomy derived from diffusion tensor imaging of the human brain.
-It incorporates multiple thalamic nuclei and a six-layered cortical microcircuitry, which is grounded in three-dimensional reconstruction of single neurons from the cat visual cortex.
-It boasts nearly half a billion synapses, incorporating accurate receptor kinetics, short-term plasticity, and dendritic spike-timing-dependent synaptic plasticity (dendritic STDP).
-The model can demonstrate behavioral patterns that mimic normal brain activity.
-This model is a comprehensive attempt to simulate the whole neural system, which is a critical component of the brain involved in sensory processing and consciousness.
-Eliasmith et al.[29] introduce a more comprehensive brain simulating model, which consists of 2,500,000 SNN neurons.
-This model receives only sequences of visual images and produces all its responses through a simulated arm.
-Despite its simplifications of real structure, the model encapsulates numerous elements of neuroanatomy, neurophysiology, and behavioral psychology, as evidenced by its performance across eight distinct tasks.
-As for the standardized platform, Zeng et al.
-[30] presents BrainCog, a cognitive intelligence engine inspired by the brain.
-This platform, based on SNN, offers critical infrastructure for the creation of brain-inspired artificial intelligence and simulation of the brain.
-BrainCog encompasses a variety of biological neurons, encoding methods, learning principles, brain regions, and a combination of hardware and software as its core elements.
-However, large scale brain simulations require the simulation of millions or even billions of neurons, posing extremely high demands on computational resources.
-Even the most powerful supercomputers might need weeks or even months to complete a single simulation.
-Additionally, as the number of neuron connections and synaptic weights increases, data storage and processing requirements also significantly rise.
-Besides, although large scale brain simulations aim to simulate neurons and their connections, they often fail to capture all the biological details present in the brain.
-Factors like neuron morphology, molecular-level dynamics, and changes in gene expression are complex biological phenomena that are difficult to fully integrate into the models.
-Thus, the simulation results might not fully reflect the real behavior of the brain.
-At last, building a large-scale brain simulation model requires a vast amount of precise biological data, such as the connectivity between neurons, synaptic strengths, types, and concentrations of neurotransmitters.
-However, acquiring these data is often very challenging, involving complex and expensive experimental techniques, and it’s difficult to measure comprehensively across the entire brain.
-As for the collective action of neuron populations, despite the success of neural system mean field models[31] in areas such as modeling epileptic seizures[32], there is still a lack of models that precisely describe the large-scale activity of whole brain.
-A.2 Brain Network Communication Model
-The network communication model is a conceptual framework which describes how data is transmitted from one node to another over a network.
-It breaks down the complex process of communication into simpler, more manageable layers or components.
-The most widely recognized network communication models are the Open Systems Interconnection (OSI) model and the Internet Protocol Suite, commonly known as the TCP/IP model.
-Nowadays, scientists apply network communication models to neuron research, which are mostly used to understand how brain networks operate[33, 34, 35].
-Early studies have indicated that, similar to social and electrical networks[36], brain networks also exhibit properties such as power-law distributions[37] and the rich-club phenomenon[38].
-These results indicate that brain networks contain a wealth of network statistical characteristics.
-Researchers designed many brain network models to describe the processes of message propagation in brain.
-One category of brain network communication models is the routing protocol model.
-The most famous model of this category is shortest path model[39], which assumes that signals between a pair of nodes travel along the shortest path between them.
-In the shortest path model, the shortest path length from the starting node $s$ to each node $v$, is initialized as $d(s)=0$ and $d(v)=\infty$; then, process each node in increasing order of $d(v)$, for each node, update the $d(u)$ value of its neighbors $u$.
-The update way can be represented as follows:
-$d(u)=\min(d(u),d(v)+w(v,u)),$ (9)
-where $w(v,u)$ is the weight of edge $e_{uv}$.
-Information is propagated entirely through the shortest paths in these models, which means in terms of propagation cost, signal transmission is optimal.
-However, the key disadvantage of the model is these methods require each neuron must have knowledge about the entire network structure[40], which doesn’t correspond to the reality: every neuron only has the message of its neighbor neurons.
-Another well-known model is navigation model.
-It suggests that neurons always send information to the neuron that can be reached with the least communication cost[41].
-These models achieve the lowest communication cost.
-A significant achievement is that some algorithms can identify paths with 70%-100% of optimal signaling efficiency in the functional connectomes of different organisms[42].
-However, the main disadvantage of navigation model is it supposes that neurons know the distance between their neighbors and a desired target region.
-Recently their still has a lack of evidences that reveal how this knowledge would be given to neurons in brain networks.
-Another well-known category of models is the diffusion model.
-The diffusion model posits that signals propagate through random walk dynamics.
-This process of propagation does not require individual neurons to possess any prior knowledge beyond their immediate neighboring neurons.
-Among the most classic of these is the random walk model[43].
-In the random walk model, signals are transmitted to randomly selected neighboring nodes with a probability proportional to the connection weights until they reach the target node.
-For unbiased random walk, the function of random walk model is as follows.
-$P_{ij}=\frac{1}{k_{i}},$ (10)
-where $P_{ij}$ represent the probability of node $i$ choosing node $j$.$k_{i}$is the degree of node $i$.
-As for biased random walk, for example, if the transition probability depends on the edge weights or certain properties of the nodes, then the transition probability can be adjusted to
-$P_{ij}=\frac{w_{ij}}{\sum_{l\in N(i)}w_{il}},$ (11)
-where $w_{ij}$ is the weight of node $i$ to node $j$,$N(i)$is the set of node $i$.
-The diffusion process does not require individual neurons to have knowledge about the connectivity of connectome beyond their immediate neighbors.
-However, studies have shown that the efficiency of signal transmission through random walks is relatively low[44].
-So it can’t describe the rapid communication process in the brain.
-An effective category is threshold model[45].
-The threshold model was originally proposed to characterize collective behavior.
-In the threshold model, a node’s behavior depends on the behavior of its neighboring nodes.
-Suppose that within a group, nodes can hold one of two states regarding a certain behavior: an active state and a quiescent state.
-For a node, it will only transition from the quiescent state to the active state if the stimulus given by its neighboring nodes in the active state surpasses a certain threshold; otherwise, it remains in the quiescent state.
-The condition for node $i$ to adopt the behavior at time $t$+1 can be formulated as follows.
-$a_{i}(t+1)=1\quad\text{if}\quad\frac{1}{k_{i}}\sum_{j\in N_{i}}a_{j}(t)\geq\theta_{i},$ (12)
-where $N_{i}$ is the set of neighbors of node $i$, $a_{i}(t)$ is the adoption state of node $i$ at time $t$.
-$\theta_{i}$ is the threshold of node $i$.
-$k_{i}$ is the degree of node $i$.
-The threshold model has stronger biological interpretability because it approximates the mechanism of neuronal firing.
-Consequently, the threshold model has been extensively applied in artificial neural networks.
-However, how to set the propagation thresholds for different neurons, and how to design the dynamics model between neurons remains a challenge.
-Appendix B Data and Methods
-B.1 Drosophila Connectome Data Set
-We employ the recently released largest Drosophila brain connectome[17], offering a thorough mapping of the entire neural network within a Drosophila female brain.
-This data set encompasses over 120,000 neurons and 30,000,000 synapses, including detailed information on synaptic connections between neurons.
-Notably, it provides high-precision 3D coordinates, accurate to the nanometer level, for both cell bodies and synapses.
-Additionally, this data set offers extensive labeling for various types of neurons.
-For a more in-depth understanding, please refer to [17][46] for detailed information.
-In this section, we use nodes to represent neurons and connections to represent the synapses between neurons.
-Table 3 presents the basic properties of the data set.
-Note that we combine multiple edges between two nodes to calculate network statistic properties.
-The average degree before and after combining multiple edges is 501.6 and 37.01, respectively, indicating it is prevalent that the pairs of neurons are connected by multiple synapses.
+Appendix A Data and Methods
+A.1 Drosophila Connectome Data Set
+We utilize the recently released largest Drosophila brain connectome[12], which provides a comprehensive mapping of the entire neural network within a female Drosophila brain.
+This data set includes over 120,000 neurons and 30,000,000 synapses, featuring detailed synaptic connections and high-precision 3D coordinates (accurate to the nanometer level) for both cell bodies and synapses.
+Additionally, it offers extensive labeling for various neuron types.
+For more detailed information, refer to [12, 16].
+In this section, neurons are represented as nodes and synapses as connections between them.
+Tab.
+2 presents the basic properties of the data set.
+We combine multiple edges between two nodes to calculate network statistics.
+The average degree before and after combining multiple edges is 501.6 and 37.01, respectively, indicating the prevalence of multiple synapses between pairs of neurons.
 Table 2: Statistical Properties of Drosophila Brain Connectcome Graph
@@
 $501.6$ $0.1527$ $0.0008817$
-Network Diameter Assortativity Coefficient
-$11$ $0.0008817$
-B.2 Model
+Network Diameter Average Shortest Path Assortativity Coefficient
+$11$ $3.9061$ $-0.05868$
+A.2 Model
 In this section, we introduce the large scale network communication model which consumes less computational resource and can efficiently utilize the network structure.
@@
 In addition, we replace the node model in the network communication model with an improved LIF neuron model for comparison.
-B.2.1 Symbol Systems
-The original network presentation of neural systems often uses nodes to represent the neurons and using edges to represent the links (chemical synapses, electrical synapses, etc) between neurons[4].
-However, this approach has a number of limitations: first, there are usually more than one link (synapse) between two neurons[49] and the links are spatially and temporally asymmetric.
-To illustrate, two neurons shown in Fig.S1 have multiple synapses between them, and the dynamical properties of these different synapses may vary.
-Additionally, due to their spatial locations, the signal transmission delay and the position of action on the downstream neurons are also different, making it an oversimplification if the synapses are merged into one edge.
+A.2.1 Symbol Systems
+The original network representation of neural systems often uses nodes to represent neurons and edges to represent the links (chemical synapses, electrical synapses, etc.) between neurons[3].
+However, this approach has several limitations: First, multiple synapses typically exist between two neurons[17], and these synapses are spatially and temporally asymmetric.
+For instance, two neurons, as illustrated in Fig.
+S1, can have multiple synapses with varying dynamic properties.
+Additionally, due to differing spatial locations, signal transmission delays and the positions of action on downstream neurons also vary.
+Merging these synapses into a single edge oversimplifies the complexity of neural communication.
 Figure S1: An example of complex connection
@@
 $Y_{j}=1$ if the set $e_{j}$ is active, otherwise $Y_{j}=0$.
-B.2.2 Large Scale Network Communication Model
-By modeling the Drosophila connectome as graph model, we introduce the large scale network communication model on this graph.
-We use a linear threshold activation function on each neuron.
-Thus, general transmission function is defined in the following manner,
+A.2.2 Large Scale Network Communication Model (LSNC)
+By modeling the Drosophila connectome as graph, we introduce the large scale network communication model on this graph. The general transmission function is defined in the following manner,
 $Y_{j}=\sigma(\frac{1}{\beta_{j}}\sum_{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset}\sum_{k|n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\omega(n_{k},e_{i},e_{j})X_{k}),$ (3)
@@
 $\beta_{j}$ is the normalization factor.
-In this paper we use threshold activation function and sigmoid function, which are two .
+In this paper we use threshold activation function and sigmoid function for comparison.
 The threshold activation function is a simple, binary decision function used in artificial neurons.
@@
 $Y_{j}=\sigma(\frac{1}{\beta_{j}}\sum_{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset}\sum_{k|n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\omega(n_{k},e_{i},e_{j})X_{k}).$ (10)
-B.2.3 Neuron Dynamic Model
+A.2.3 Neuron Dynamic Model
 To better reflect real-world conditions and demonstrate that the superiority of this network architecture does not solely rely on the design of the activation function, we also employed a variant of the LIF model[4] as the basic neuron model.
@@
 Other propagation mechanism is consistent with network model introduced in section B.2.2.
-B.3 Evaluation Criteria
+A.3 Evaluation Criteria
 In previous study, researchers used classic indicators such as signaling cost and activation time to evaluate the quality of the model[7][20][21].
@@
 If regions related to the certain function have a significant higher activation percentage than the average activation percentage of the whole brain, we call it a quasi-real activation pattern.
-We define the consistency indicator $C_{r}$ in region $r$ which has a certain function (visual, olfactory, etc) is defined in the following manner.
-$C_{r}=N_{r}^{act}/N_{r},$ (26)
-where $N_{r}^{act}$ represents the number of neurons activated in area $r$ at the final state, $N_{r}$ represent the number of neurons in area $r$ and $N$ represent the total number of neurons.
-Appendix C Introduction of Related Neurons
+We define the consistency indicator $C$ in a area which has a certain function (vision, olfaction, etc) is defined in the following manner.
+$C=N^{act}/N,$ (14)
+where $N^{act}$ represents the number of neurons activated in this area at the final state, $N$ represent the number of neurons in this area and $N$ represent the total number of neurons.
+Appendix B Introduction of Related Neurons
 Optic neurons, also known as retinal ganglion cells, are the first neurons in the visual pathway.
@@
 These neurons play a critical role in modulating and refining the olfactory information by providing inhibitory or excitatory input to projection neurons and other local neurons, thus shaping the olfactory responses and contributing to odor discrimination and perception.
+Tab.
+3 represents the numbers of specific neurons of each hemisphere in related areas.
+It presents that the number of neurons in specific areas of the Drosophila brain connectome is mirror-symmetrical.
+Additionally, visual neurons constitute the largest proportion among the connectome’s neurons (approximately one-twice).
+Neurons in the olfactory-related regions, though comprising a smaller proportion (for instance, the ALON region has only 8 neurons on each side), play a critical role in olfactory perception.
+Table 3: Node numbers in related areas Drosophila Brain Connectcome Graph
+total (left) total (right) optic (left) optic (right)
+$60834$ $63144$ $36131$ $37524$
+visual projection (left) visual projection (right) ALIN (left) ALIN (right)
+$3117$ $3117$ $12$ $12$
+ALLN (left) ALLN (right) ALON (left) ALON (right)
+$196$ $194$ $8$ $8$
+ALPN (left) ALPN (right)
+$310$ $314$
 For further information, please read [16].
+Appendix C Experimental Settings
+The basic methodology of the experiment is as follows.
+We apply a continuous stimulus signal (binary 0-1 signal) to a subset of input neurons (i.e., visual or olfactory neurons) in the brain.
+We then compute the signal’s propagation process within the brain and record the responses of both intermediate and output neurons.
+We give 2 types of input neurons (visual, olfactory) stimulus and record the activation patterns of 6 type of neurons: optic (visual), visual_projection (visual), ALLN (olfactory), ALIN (olfactory), ALPN (olfactory) and ALON (olfactory).
+When we give neuron a stimulus, we set the state of neuron as 1.
+The evaluation criterion involves calculating the ratio of neurons that should be activated to those that are actually activated in a specific area (see Equ.
+14 in Appendix A).
+Given that the brain functions similarly to a shallow neural network[22], we document outcomes at an iteration step size of 5.
+Since the input is constant and no randomization parameters are introduced, the results of each experiment are consistent, eliminating the need for repetition.
+For detailed information on related neurons, please read Appendix B.
+In the linear threshold model, we set $\sigma_{0}=0.8$.
+For the simulation of the LIF model, numerous hyperparameters are required.
+All the hyperparameters used are listed in the following table.
+Biological hyperparameters are set according to real biological experimental tests[23].
+The current proportion constant is the optimal parameter obtained through equidistant search.
+Table 4: List of Hyperparameters
+Hyperparameter Description Value
+$\tau$ Membrane time constant 15.0 ms
+$V_{\text{rest}}$ Resting membrane potential -65 mV
+$V_{\text{th}}$ Spike threshold -45 mV
+[…]
+$dt$ time step 0.55 ms
+$\alpha$ Current proportion const 0.03/0.045 (visual/olfactory)
+To demonstrate the importance of network structure, we add a comparative model that randomly reconnects networks.
+Specifically, we perform degree-preserving rewiring on a certain proportion of edges in the brain, which means perturbing the network structure while ensuring that the degree of each node remains unchanged.
+In Fig.
+2 we set reconnect rate $p=20\%$.
+For its sensitivity analysis, please read Appendix E.
 Appendix D Bilateral Response under Unilateral Stimulus
 In real life, the left and right hemispheres cooperate to perform complex tasks.
-We are interested in whether network models can simulate this cooperative effect to some extent.
-So we administer a unilateral stimulus to the input neurons and monitor the varying activation ratios across both hemispheres.
-More precisely, we independently simulate visual input neurons in the left and right hemispheres and record the rate of activated neurons in each hemisphere over time.
-Fig.S3 illustrates the curve depicting how the response ratio evolves with each time step.
-The horizontal line represents the proportion of activated neurons in all hemisphere.
-It is noteworthy that when simulating one hemisphere alone, visual processing neurons in both hemispheres are activated.
-At time step 5, the unstimulated hemisphere begins to have a activate pattern, which corresponds to the hypothesis of shallow connection[53].
-Furthermore, our observations reveal a concurrent effect: the activation pattern in one hemisphere mirrors that in the unstimulated hemisphere.
-These findings are reflective of reality.
-Although the left and right hemispheres of the brain possess certain functional specializations, they also demonstrate a synergistic effect.
-Typically, they collaborate to fulfill specific functions[56].
-For instance, during complex cognitive tasks, both hemispheres coordinate their efforts to contribute to task completion.
-Even when simulating input to neurons on just one side of the brain, our model reveals similar synergistic effects, indirectly attesting to the model’s effectiveness.
-Besides, the mirror effect align with real situation, underscoring the brain’s resilience as a system.
-In instances of disease or physical damage leading to impairment in one hemisphere, this interconnected structure supports the continuation of critical functions, such as vision.
+To examine whether network models can simulate this cooperation, we administered a unilateral stimulus to input neurons and monitored the varying activation ratios across both hemispheres.
+Specifically, we independently simulated visual input neurons in the left and right hemispheres and recorded the rate of activated neurons in each hemisphere over time.
+Second, we investigate whether local attack to one side of the network—analogous to real-world conditions such as tumors and other diseases—affects this bilateral response process, and how the extent of such damage influences the synergistic response.
+The specific methods of inducing attack are as follows: We first identify two symmetrical coordinates on both hemispheres of the brain.
+Using these coordinates as the center, we remove nodes within varying radii and subsequently record the response curves for each hemisphere-related region.
+Fig.
+S3 illustrates the specific length measurements of the Drosophila connectome and the regions examined in this study.
+The purple and blue cells represent neurons associated with visual processing.
+Spheres $S1$, $S2$, $S3$, and $S4$ each represent four spherical regions subjected to localized attacks.
+Sphere $S1$ and $S2$ encompass approximately 2,000 neurons each, while sphere $S3$ and $S4$ each cover approximately 12,000 neurons.
+(a) Geometric dimensions of Drosophila connectome
+(b) Locations of targeted regions
+Figure S3: Schematic diagram of the specific length measurements of the Drosophila connectome and the location of regions deleted. The units of center and r in labels are micrometers.
+Fig.
+S4 illustrates the curve depicting how the response ratio evolves with each time step.
+We use $C_{avg}$ to represent the average activated rate of areas related to vision.
+Notably, when stimulating one hemisphere alone, visual processing neurons in both hemispheres are activated (see curve regular left and right in Fig.
+S4(a) and (b)).
+[…]
+Below, we further explain the reasons for the significant delay in responses to unilateral large-scale regional attacks on this side’s stimuli (e.g., in Fig.
+S4(a), where the local damaged area is the $S2$ region of the left hemisphere, and stimuli on the left side cause responses in both the left and right hemispheres (light purple and dark purple curves) to be delayed).
+Calculations show that there are 19,031 synapses from the left hemisphere’s visual input area to the left hemisphere’s visual signal processing area (optic, visual_projection), while there are only 18 synapses from the left hemisphere’s visual input area to the right hemisphere’s visual signal processing area.
+Furthermore, there are 338,456 synapses from the left hemisphere’s visual processing area to the right hemisphere’s visual signal processing area.
+This implies that, primarily, the unilateral visual input area essentially does not interact with the contralateral visual signal processing area, and interactions occur indirectly through the ipsilateral visual processing area.
+Consequently, when the visual signal processing area of this side is damaged, the reduction in the number of neurons makes it challenging for the remaining neurons to reach an excitatory state, causing delays in the ipsilateral visual signal processing area.
+This delay, in turn, leads to delayed responses in the contralateral visual signal processing area.
 (a) Left stimulus
 (b) Right stimulus
-Figure S3: Experiment of activation rate under single stimulus. We set threshold $\sigma_{0}=0.8$.
+Figure S4: Experiment of activation rate under single stimulus.
+All responses on the left hemisphere are marked with light-colored lines, and all responses on the right hemisphere are marked with dark-colored lines.
+We set threshold $\sigma_{0}=0.95$.
 Appendix E Impact of Perturbation Rate on Activation Areas
-The main text demonstrates that with a 10% perturbation in the network, the original realistic activation patterns vanish regardless of the model used.
 In this section, we discuss the impact of different reconnection rates on activation patterns using the degree-preserving reconnection method.
-The experiments presented in this section involve calculating patterns induced by two types of stimuli across three models under network perturbations of 1%, 3%, 5%, 10%, and 20% reconnect rates.
-We normalized each data point using the difference between the activation ratio of region $r$ and the average activation ratio $\bar{C}$ of the entire brain, that is to say, $\Delta C=C_{r}-\bar{C}$.
-The experimental results are shown in Fig.S4, where 0 represents the original, unperturbed experimental results.
-It can be observed that at the 0 point, there is a significant difference in the relative activation ratios between visual and olfactory areas under specific stimulus.
-However, this significant realistic activation patterns which we discussed in the main text disappear with a 1% network perturbation, replaced by a roughly equal activation ratio across all regions.
-This indicates that even a 1% change in network structure can lead to a rapid loss of network functionality specificity.
-Additionally, it is noted that small-scale network perturbations (1%-5%) have a lesser impact on olfactory stimulus compared to visual stimulus.
+The experiments presented in this section involve calculating patterns induced by two types of stimuli across three models under network perturbations of 1‰, 5‰, 1%, 3%, 5%, 10%, and 20% reconnect rates.
+Let $C$ be the average activation ratio of an area.
+We normalized each data point using the difference between the activation ratio of this area and the average activation ratio $\bar{C}$ of the entire brain, that is to say, $\Delta C=C-\bar{C}$.
+The experimental results are shown in Fig.
+S5, where 0 represents the original, unperturbed experimental results.
+At the 0 point, there is a significant difference in the relative activation ratios between visual and olfactory areas under specific stimuli.
+However, this significant activation pattern, discussed in the main text, disappears with a 1‰’s network perturbation, replaced by a roughly equal activation ratio across all regions.
+This indicates that even a 1‰’s change in network structure can lead to a rapid loss of network functionality specificity.
+Additionally, small-scale network perturbations (1‰-5%) have a lesser impact on olfactory stimuli compared to visual stimuli.
 This is because the number of neurons associated with olfactory functions is significantly fewer than those related to visual functions, making it less likely for small-scale network structure changes to affect these neurons, thereby leading to a smaller impact on functionality.
-At the same time, various horizontal comparisons indicate that while there are no significant differences among different models in forming activation patterns, more complex neuron models have a stronger ability to maintain their original activation patterns under network disturbances.
-This is because the signals generated by complex neuron models have greater redundancy and functional diversity.
+Horizontal comparisons indicate that while there are no significant differences among different models in forming activation patterns, more complex neuron models better maintain their original activation patterns under network disturbances.
+This is because signals generated by complex neuron models have greater redundancy and functional diversity.
 For instance, the LIF (Leaky Integrate-and-Fire) model has a voltage decay mechanism, where the voltage gradually decays in the absence of a signal.
 Even if some input signals are erroneous, as long as the proportion of errors is not very large, the decay mechanism can filter out the impact of these erroneous signals.
-These results indicate that the formation of realistic activation patterns is highly sensitive to changes in network structure, with even minor perturbations capable of disrupting this network configuration.
-This indirectly substantiates that the authentic network structure is crucial for the formation of genuine activation patterns.
-Additionally, it highlights that complex functions (such as vision) are more susceptible to network disturbances.
 (a) threshold model, visual stimulus
@@
 Appendix F Large Spatial Network Visualization
-The visualization software of experiment observation on large spatial networks is important for researchers[57][58].
+The visualization software of experiment observation on large spatial networks is important for researchers.
 However, there is still a lack of visualization system that can observe information propagation on large spatial networks in real-time, which poses difficulties in experiment.
@@
 F.3 Function Display
-Fig.SS5 is an example of the 3D visualization system.
+Fig. S6 is an example of the 3D visualization system.
 (a) Side view 1
@@
 (i) Front view 3
-Figure S5: Visualization of Spatial Drosophila Brain Network
-All nodes’ colors can be adjusted by visible selection or input RGB code (Fig.SS6).
-Figure S6: Display of color selection
-Besides, from square to round, the shape of nodes can be adjusted even real neuron (Fig.SS7).
-Figure S7: Display of color selection
-It also can present different single areas (Fig.SS8).
-Figure S8: Display of color selection
+Figure S6: Visualization of spatial Drosophila brain network.
+All nodes’ colors can be adjusted by visible selection or input RGB code (Fig. S7).
+Figure S7: Display of color selection.
+Besides, from square to round, the shape of nodes can be adjusted even real neuron (Fig. S8).
+Figure S8: Display of nodes size adjusting.
+It also can present different single areas (Fig. S9).
+Figure S9: Display of area selection.